### v1

# ITDR: An Instruction Tuning Dataset for Enhancing Large Language Models in Recommendations

Abstract.
Large language models (LLMs) have demonstrated outstanding performance in natural language processing tasks. However, in the field of recommendation systems, due to the structural differences between user behavior data and natural language, LLMs struggle to effectively model the associations between user preferences and items. Although prompt-based methods can generate recommendation results, their inadequate understanding of recommendation tasks leads to constrained performance. To address this gap, in this work, we construct a sufficient instruction tuning dataset, ITDR, which encompasses 7 subtasks across two core root tasks—user-item interaction and user-item understanding. The dataset integrates data from 13 public recommendation datasets and is built using manually crafted standardized templates, comprising approximately 200,000 instances. Experimental results demonstrate that ITDR significantly enhances the performance of mainstream open-source LLMs such as GLM-4, Qwen2.5, Qwen2.5-Instruct and LLaMA-3.2 on recommendation tasks. Furthermore, we analyze the correlations between tasks and explore the impact of task descriptions and data scale on instruction tuning effectiveness. Finally, we perform comparative experiments against closed-source LLMs with substantial parameters. Our tuning dataset ITDR and the fine-tuned large recommendation models can be accessed at https://github.com/hellolzk/ITDR.
Recommendation system, Large language model, Instruction tuning, Datasets
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2026/08ccs: Information systems Recommender systems
Figure 1. ITDR architecture: tasks, subtasks, and datasets.
1. Introduction
Recommendation systems, as a core technology to address the issue of information overload, have been driving the innovation of personalized user experiences in scenarios such as e-commerce, social media, and content platforms (REVIEWOFRECOMMENDERSYSTEMS). With the groundbreaking advancements in LLMs (zhao2025surveylargelanguagemodels), how to effectively leverage LLMs to enhance the performance of recommendation systems has become a cutting-edge topic of shared interest in both academia and industry (lin2025can). Among these efforts, instruction tuning techniques are regarded as a critical method to achieve this goal (RecLM; tallrec; Recommendationasinstructionfollowing; Parameter-efficientconversationalrecommender; lin2024rella; Data-efficientFine-tuningforLLM-basedRecommendation), owing to their ability to significantly improve the model’s understanding of specific user intents and task requirements.
However, current recommendation methods that rely on instruction-tuning of LLMs face a serious data bottleneck problem (wu2024surveylargelanguagemodels). First, traditional recommendation datasets primarily consist of ID-based explicit or implicit user-item interaction records (e.g., ratings or clicks), generally lacking structured task descriptions and diverse natural language instructions—precisely the prerequisites for effective fine-tuning of LLMs on recommendation tasks. Second, the diverse nature of recommendation tasks demands that training data comprehensively cover different types of task scenarios and user behavior patterns. The limitations of existing datasets severely constrain the generalization and adaptability of LLMs in recommendation scenarios, highlighting the urgency of constructing high-quality, diverse instruction datasets.
To address these challenges, we propose ITDR—a large-scale Instruction Tuning Dataset specifically tailored for Recommendation systems. ITDR integrates multi-dimensional data from 13 classic recommendation benchmarks (including MovieLens (MovieLens), Amazon Reviews (AmazonReviews2023), etc.), categorized into two core task classes—user-item interaction (UII) and user-item understanding (UIU)—and further subdivided into seven distinct subtasks. Each subtask is equipped with a meticulously designed task description to guide the LLM’s decision-making behavior. These subtasks comprehensively cover key recommendation scenarios such as rating prediction, sequential recommendation, and interest recognition, ensuring the dataset’s broad applicability. ITDR ultimately comprises nearly 200,000 high-quality instructions, and we commit to publicly releasing all datasets and fine-tuned models to foster continued progress in this field.
Based on the ITDR, we conducted systematic instruction tuning experiments on multiple open-source LLMs. Experimental results demonstrate that ITDR significantly enhances the performance of language models of varying scales across diverse recommendation tasks.
In summary, the key contributions of this work are as follows:
1. The construction and open-sourcing of the ITDR benchmark dataset—a large-scale, task-comprehensive instruction fine-tuning dataset for recommendation systems, containing nearly 200,000 high-quality instructions across two root tasks and seven subtasks;
2. Rigorous experimental validation of ITDR’s effectiveness in enhancing the recommendation performance of LLMs, with thorough ablation studies providing critical references for future optimization and dataset expansion in recommendation-oriented LLMs.
2. Related Work
2.1. Large Language Models for Recommendation
LLMs have demonstrated remarkable capabilities in natural language understanding and generation, achieving successful applications in various tasks such as text summarization and question answering. Leveraging their strengths in capturing complex linguistic patterns, researchers have begun exploring the potential of LLMs in recommender systems. Current research can be broadly categorized into two main approaches.The first approach directly employs LLMs as recommender systems by designing sophisticated prompt templates to harness their zero/few-shot (few-shot) learning capabilities. Representative works include: Chat-REC (Chat-rec), which transforms user profiles and interaction histories into prompts to enable conversational recommender systems; LLM4CDR (LLM4CRD) constructs context-aware prompts based on user’s purchase histories in source domains and cross-domain shared features to enhance cross-domain recommendation performance; and RecPrompt (Recprompt) develops an adaptive prompt framework for news recommendation that automatically optimizes prompt design.The second approach integrates LLMs with traditional recommender systems to augment their capabilities. Notable examples include Llmrec (Llmrec), which enhances graph representations of user-item interactions by incorporating LLM-generated features to alleviate data sparsity in implicit feedback scenarios, and LLM-CF (LLM-CF) distills the world knowledge and reasoning capabilities of LLMs into collaborative filtering frameworks, significantly improving recommendation effectiveness. While these methods can enhance LLMs‘ capabilities in recommendation tasks to some extent, they essentially rely on prompt engineering to elicit emergent abilities and align the models with recommendation objectives. However, due to the complexity of user behavior patterns and item features, as well as the scarcity of recommendation-related concepts in natural language text, larger models without fine-tuning do not perform well in some recommendation tasks (e.g., sequence recommendation) (Ischatgptagoodrecommender).
Unlike existing studies, our work focuses on systematically enhancing the recommendation capabilities of LLMs through instruction tuning. By improving the model’s semantic understanding of recommendation instructions and task responsiveness, we significantly enhance its practicality and adaptability in diverse recommendation scenarios.
2.2. Instruction Tuning for LLMs
Instruction tuning is a methodology for adapting pre-trained LLMs using structured datasets expressed in natural language (Finetunedlanguagemodelsarezero-shotlearners). This training paradigm enhances a model’s ability to comprehend and execute human instructions. Conceptually, IT shares similarities with both supervised fine-tuning approaches (Traininglanguagemodelstofollowinstructionswithhumanfeedback) and multi-task learning frameworks based on prompt-based training (Multitaskpromptedtrainingenableszero-shottaskgeneralization). The effectiveness of instruction tuning arises from its dual capacity: it not only improves performance on tasks encountered during explicit training but also fosters generalization capabilities that transfer to novel, unseen tasks (Multitaskpromptedtrainingenableszero-shottaskgeneralization).
In this work, we focuse on optimizing the performance of LLMs in recommendation system-related tasks through instruction tuning techniques. We construct a meticulously designed dataset that fully incorporates the unique characteristics and practical application scenarios of recommendation tasks. Experimental results show that the adopted instruction tuning method can significantly enhance the model’s comprehensive performance across various recommendation tasks, validating the effectiveness and applicability of this technology in the field of recommendation systems.
3. ITDR: Instruction Tuning Dataset for Recommendation
In this section, we provide a detailed introduction to the ITDR dataset designed for instruction tuning tasks in recommender systems. Given the substantial resources required to construct a comprehensive instruction tuning dataset that covers diverse recommendation scenarios, we adopt the methodology of prior work (INTERS) by transforming existing recommendation datasets into an instructional format for efficient construction. Based on the characteristics of recommendation tasks, we categorize all tasks and their corresponding datasets into two root tasks: User-Item Interaction (UII) and User-Item Understanding (UIU). The complete taxonomy of tasks and their associated datasets is illustrated in Figure 1. Below, we elaborate on the specific definitions of these two root tasks, and the data construction process.
3.1. Root Task Definition
3.1.1. User-Item Interaction(UII)
In recommendation systems, user-item interactions refer to behavioral records between users and the content recommended by the system, typically including explicit or implicit signals such as clicks, views, purchases, and ratings. The core objective is to characterize the relationship between user preferences and item features, thereby providing a basis for personalized recommendations. In the recommendation process, modeling user-item interactions is a critical factor determining the accuracy of the recommendation system and the user experience. To this end, we employed four classical interaction recommendation tasks: Rating Prediction(RP), Top-K Recommendation(TKR), Cross-Domain Recommendation(CDR) and Next Item Recommendation(NIR), requiring the model to mine latent patterns in the interaction data and predict user’s interest levels in unexposed items.
3.1.2. User-Item Understanding(UIU)
In recommendation systems, user-item understanding refers to the in-depth feature mining from the item dimension. This process involves analyzing the inherent attributes of items (such as multi-dimensional features of movies, including genre, director, cast, plot synopsis, etc.) to construct accurate item profiles, thereby achieving personalized recommendation matching for target users. We design three tasks: User Attribute Prediction(UAP), Interest Recognition(IR) and Target User Identification(TUI), requiring the model to deeply comprehend item features and demonstrate better adaptability in personalized recommendation scenarios.
3.2. ITDR Construction
With the task defined and dataset selected, we construct ITDR through the following three steps.
1. Raw data collection and preprocessing. We manually collect various datasets related to recommender system tasks from multiple authoritative open data sources. These datasets generally exhibit high quality assurance for two key reasons: (1) they are widely adopted as benchmark datasets in the field of recommendation; (2) their data quality has been empirically validated through extensive prior research. This dual verification mechanism ensures the reliability and academic value of the collected data. After collection, we removed irrelevant item features and disqualified users.
2. Constructing template and task descriptions. We design customized templates for each dataset, which are closely aligned with specific task requirements through natural language instructions. Additionally, we craft detailed task descriptions for each subtask to more effectively guide the model in accomplishing recommendation tasks.
3. Example generation We use the task description as the ”instruction” part of the example and the template as the ”input” part of the example. For the UII task, the ground truth corresponding to each instruction can be directly obtained from the original dataset. As for the ”target user identification” and ”interest recognition” subtasks in the UIU task, since the original dataset does not include the corresponding true annotation data, we employ DeepSeek-V3 (Deepseek-v3) to generate the reference ground truth for these tasks.
After completing the above steps, we construct an ITDR instruction set comprising 2 root tasks and 7 subtasks, with a total scale of 195,065 instructions. Appendix A provides a detailed description of the root tasks and subtasks. Appendix B compiles comprehensive statistics and complete experimental results for ITDR. Appendix C offers specific illustrative examples for each subtask.
4. Experiments
In this section, we conducted extensive experiments in order to address the following key research questions:
$\bullet$ RQ1: Can instruction fine-tuning with ITDR enhance models’ recommendation capabilities?
$\bullet$ RQ2: Do potential interactions exist between different root tasks?
$\bullet$ RQ3: What are the underlying relationships among subtasks?
$\bullet$ RQ4: To what extent do our designed task descriptions guide models in recommendation tasks?
$\bullet$ RQ5: How does data scale affect model performance?
$\bullet$ RQ6: What are the performance differences between fine-tuned models and closed-source LLMs on recommendation tasks?
4.1. Backbone Models
We select four publicly available LLMs spanning different scales: GLM-4-9B (glm4), Qwen2.5-7B (qwen2.5) and its instruction-tuned variant Qwen2.5-7B-Instruct (qwen2.5), and LLaMA-3.2-3B (llama3.2). These models represent diverse architectures that have demonstrated strong performance on standard benchmarks and are widely applied in a variety of research studies.
4.2. Experiment Setup
4.2.1. Dataset Split
ITDR comprises two root tasks with seven subtasks, totaling 195,065 data instances. In order to balance computational efficiency and fine-tuning performance, we extracted 39,061 instances and partitioned them into 31,250 training instances and 7,811 testing instances following proportional stratified sampling. All backbone models were fine-tuned and evaluated based on these data.
4.2.2. Evaluation Metrics
For numerical prediction tasks, including rating prediction and the prediction of user attributes such as age and birth year, we use Root Mean Squared Error (RMSE) (RMSE) as the evaluation metric. For other user attribute predictions, such as gender, location, and occupation, whose outputs are discrete categories, we use Accuracy (ACC) to evaluate the model’s performance. Our assessment of top-k recommendation and cross-domain recommendation tasks employs metrics widely adopted in prior research (NDCG): N@10 (Normalized Discounted Cumulative Gain at 10), H@1 (Hit Ratio at 1), and P@5 (Precision at 5). In next-item prediction, we exclusively report the H@1 result, focusing on the model’s ability to accurately predict the user’s next interacting item. Finally, for interest recognition and target user identification, which are inherently generative tasks, we evaluate the models’ generation capabilities using BLEU-1&2 and ROUGE-1&2&L (bleu; rouge). These metrics are standard for assessing the quality of generated text by comparing it against reference outputs.
Furthermore, building upon the methodology established in (INTERS), we define two composite metrics to holistically evaluate model performance on both user-item interaction and user-item understanding tasks. The construction methodology is as follows:
We first define the set of subtasks (ST) and the set of evaluation metrics for sequential recommendation tasks (M) as follows:
$ST=\{RT,TKR,CDR,NIR,UAP,IR,TUI\},$
$M=\{N@10,H@1,P@5\}.$
For each subtask, all the data it contains can be represented as the following set:
$D_{st}=\left\{D_{st,1},D_{st,2},\cdots,D_{st,n}\right\},$
where $st\in ST$, $D_{st,n}$ represents an arbitrary record in the subtask, the number of data entries contained in each subtask is $\left|D_{st}\right|$. The average performance of user-item interaction is defined as follows:
(1) $pUII=\frac{1}{8}$ $\left[\sum_{D_{j}\in\left\{D_{TKR},D_{CDR}\right\}}\frac{\sum_{m\in M}\sum_{i\in D_{j}}m(i)}{\left|D_{j}\right|}\right.$
$+\frac{\sum_{m\in M\backslash\{N@10,P@5\}}\sum_{i\in M}m(i)}{\left|D_{NIR}\right|}$
[…]
For the UAP subtask, since the evaluation metrics vary across different datasets, we define:
$D_{UAP}=\left\{D_{1},D_{2},D_{3},D_{4},D_{5}\right\},$
where $D_{1}$ to $D_{5}$ represent Last.FM 1K, Last.FM 360K, Anime Dataset 2023, BookCrossing, and MovieLens 1M respectively. In addition, we define the metric sets for evaluating subtasks IR and TUI as:
$M^{\prime}=\{BLEU-1\&2,ROUGE-1\&2\&L\}.$
Consequently, the average performance of user-item understanding is obtained as:
(2) $pUIU=\frac{1}{16}\left[\frac{\sum_{i\in\left\{D_{1},D_{2}\right\}}\mathrm{ACC}_{\text{country }}(i)}{\left|D_{1}\right|+\left|D_{2}\right|}+\frac{\sum_{i\in\left\{D_{\mathrm{UAP}}\backslash D_{4}\right\}}\mathrm{ACC}_{\text{gender }}(i)}{\left|D_{\mathrm{UAP}}\backslash D_{4}\right|}\right.$
$+\frac{\sum_{i\in\left\{D_{3},D_{4}\right\}}\mathrm{ACC}_{\text{location }}(i)}{\left|D_{3}\right|+\left|D_{4}\right|}+\frac{\sum_{f\in\{\text{ occupation,age group }\}}\sum_{i\in\left\{D_{5}\right\}}\mathrm{ACC}_{f}(i)}{\left|D_{5}\right|}$
$-\frac{1}{4}\left(\sum_{D_{k}\in\left\{D_{1},D_{2},D_{4}\right\}}\lg\frac{\sum_{i\in\left\{D_{k}\right\}}RMSE_{\text{age }}(i)}{\left|D_{k}\right|}+\lg\frac{\sum_{i\in\left\{D_{3}\right\}}RMSE_{\text{birth year }}(i)}{\left|D_{3}\right|}\right)$
[…]
Specifically, we apply a logarithmic normalization to the RMSE metric to ensure its scale aligns with that of the other metrics.
4.2.3. Implementation Details
To conduct our experiments, we performed LoRA fine-tuning using the LLaMA-Factory (zheng2024llamafactory) framework with BF16 mixed-precision training, setting the learning rate to 1e-4. We employed a cosine learning rate scheduler over 2 epochs, with a per-device batch size of 1 and gradient accumulation steps of 8. All experiments were executed on a cluster of 8 NVIDIA RTX 3090 24GB GPUs. For the testing phase, models were deployed and inferred using the vLLM (vllm) framework. As the primary objective of this study is to validate the effectiveness of ITDR, we excluded hyperparameter analysis from our experimental scope.
4.3. Results and Analysis of Key Research Questions
RQ1: Can instruction fine-tuning with ITDR enhance models’ recommendation capabilities?
We first fine-tune the model on the entire training set and then evaluate their performance on the test set to validate the effectiveness of using ITDR for instruction tuning in recommendation tasks.
The experimental results are presented in Figure 2. Overall, after fine-tuning on ITDR, models of varying scales demonstrate significant performance improvements, confirming the effectiveness and general applicability of instruction tuning in enhancing LLM-based recommendation capabilities. Additionally, we observe the following phenomena:
Figure 2. pUII and pUIU of all backbone models on two tasks before and after fine-tuning. ”ITDR-*” refers to the backbone model ”*” is fine-tuned by ITDR.
$\bullet$ Models with larger parameters show significant performance advantages. For instance, compared with LLaMA-3.2-3B, the Qwen2.5-7B series models and the GLM-4-9B perform better, indicating that larger parameters have inherent advantages in both tasks.
$\bullet$ In the UII task, GLM-4-9B exhibits optimal performance both before and after fine-tuning, which is attributed to its larger parameter count compared to other models. Notably, the experimental results show that the performance of the Qwen2.5-7B-Instruct without fine-tuning is lower than that of Qwen2.5-7B, suggesting that the instruction-tuned version may not outperform the base version in the UII task scenario.
$\bullet$ In the UIU task, the Qwen2.5-7B-Instruct without fine-tuning exhibits the best performance among the four models. This phenomenon can be attributed to the model’s superior instruction-following capability, which effectively bridges the gap in parameter magnitude between it and the GLM-4-9B model. However, ITDR-GLM-4-9B achieves the optimal performance, a result that further validates the significant advantages of large-scale parametric models in terms of task learning and generalization capabilities.
$\bullet$ After fine-tuning with ITDR, the performance disparities among the four models were significantly reduced. Taking the UII task as an example, the pUII metric gap between the un-tuned LLaMA-3.2-3B and Qwen2.5-7B exceeded 10, while after fine-tuning, this gap was reduced to within 2. These empirical results demonstrate that ITDR exhibits broad model adaptability and can effectively enhance the performance of different types of LLMs on recommendation tasks.
RQ2: Do potential interactions exist between different root tasks?
Table 1. Average performance of removing different subtasks. $\downarrow$:lower is better. $\uparrow$:higher is better. “FT” means “fine-tuning”. Bold: performance exceeds unfine-tuned model after subtask removal; underline: performance drops below unfine-tuned model.
Subtask w/o FT FT w/o RP w/o TKR w/o CDR w/o NIR w/o UAP w/o IR w/o TUI
[…]
In this scenario, we design category-level ablation studies to explore the correlation between the two root tasks. For each of the four models, the UIU and UII datasets are removed from the training set, and the rest of the data is retained for fine-tuning and then evaluated on the full test set (detailed performance metrics are shown in Figure 3). The experimental design aims to reveal the potential transfer learning mechanisms and performance associations between different root tasks.
Our experimental findings can be summarized as follows:
$\bullet$ As shown in Figure 3(a), models fine-tuned solely on the UIU dataset still demonstrate superior performance on the UII task compared to unfine-tuned models, a phenomenon consistently observed across all four models. This indicates that the semantic understanding capability acquired from the UIU data has produced a cross-task transfer effect on the UII task. Notably, when fine-tuned exclusively on UII data, the comprehensive performance of GLM-4-9B, Qwen2.5-7B, and LLaMA-3.2-3B models did not reach the level achieved by full-data fine-tuning, while Qwen2.5-7B-Instruct exhibited a unique reverse enhancement trend. We speculate that this discrepancy may stem from Qwen2.5-7B-Instruct’s specialized instruction optimization during pre-training, which enhances its adaptability to single-task data.
$\bullet$ Figure 3(b) presents the performance comparison of category-level ablation experiments on the UIU task. The result shows that for GLM-4-9B, Qwen2.5-7B, and LLaMA-3.2-3B, fine-tuning with only UIU data consistently yielded slightly inferior results compared to the full-data fine-tuning approach. However, Qwen2.5-7B-Instruct once again exhibits unique characteristics, with its single-task fine-tuning performance surpassing that of full-data training. This phenomenon further validates the heightened sensitivity of instruction-optimized models to single-task data. Particularly noteworthy is that when fine-tuned solely on UII data, all models showed significant degradation in performance on the UIU task, even falling below the level of unfine-tuned models. We attribute this performance decline to the fine-tuning process causing the model parameters to excessively favor UII task-specific feature representations, thereby compromising the general semantic understanding capabilities acquired during pre-training. The underlying mechanisms of this phenomenon remain to be further investigated.
(a) Average performance of UII.
(b) Average performance of UIU.
Figure 3. Average performance of removing different root tasks.
RQ3: What are the underlying relationships among subtasks?
In this section, we investigate the synergistic mechanisms among subtasks through task-level ablation experiments. Specifically, we sequentially remove specific subtasks (e.g., the Top-k recommendation task) from ITDR, then fine-tune GLM-4-9B using the remaining training data, and evaluate its performance on the complete test set.
The experimental results are shown in Table 1. The key experimental results are summarized below:
$\bullet$ The model demonstrates task-level generalization capabilities. For instance, after removing fine-tuning for CDR and NIR tasks, the model’s performance on these tasks still surpasses that of the unfine-tuned model. We posit that the knowledge acquired through TKR task can be effectively transferred to other sequential recommendation tasks, likely because these tasks fundamentally represent different variants of sequential recommendation tasks. Similarly, the same generalization pattern is observed in IR task: despite lacking dedicated fine-tuning for IR task, the model can still leverage knowledge learned from other tasks to benefit IR task’s performance. These findings confirm the model’s ability for cross-task knowledge transfer and generalization.
$\bullet$ The knowledge transfer between different subtasks does not always yield positive effects. Taking RP, TKR, and TUI tasks as examples, when their training data is excluded, the model’s performance actually deteriorates compared to the unfine-tuned control group. This phenomenon not only reveals the complex interdependencies among different task datasets, but also indicates that simple task combinations may lead to negative knowledge transfer effects. These findings highlight the importance of in-depth research on data composition optimization in instruction tuning, suggesting the need to develop more rational data selection strategies to achieve optimal synergistic effects between tasks.
Table 2. Ablation studies of subtasks in removing task descriptions during fine-tuning. $\downarrow$:lower is better. $\uparrow$:higher is better. “FT” means “fine-tuning”. The best performance is denoted in bold.
Subtask w/o FT w/o Description FT
[…]
(a) Ablation studies of root tasks in removing task descriptions during fine-tuning.
(b) Average performance of using different data volumes for fine-tuning.
Figure 4. Performance of using task description and different data volumes.
Figure 5. pUII and pUIU comparison of fine-tuned models with closed-source LLMs.
RQ4: To what extent do our designed task descriptions guide models in recommendation tasks?
ITDR provides detailed task descriptions for each subtask, a design intended to enhance the model’s task comprehension and enrich data diversity. In this section, we focus on evaluating the practical impact of these task descriptions on model performance. By removing the task description component from the instructions and fine-tuning the GLM-4-9B model, we obtained the experimental results illustrated in Figure 4(a) and Table 2. Based on these findings, we draw the following conclusions:
$\bullet$ Task descriptions play a critical role in the model fine-tuning process. Both pUII and pUIU exhibit a declining trend when task descriptions are removed. As shown in Table 2, the introduction of task descriptions improves model performance in most subtasks, validating the significant value of task descriptions for model training.
$\bullet$ It is noteworthy that in IR and TUI tasks, the use of task descriptions leads to a slight average performance decline. This may stem from the nature of such generative tasks: descriptive text not only fails to provide effective guidance but may also introduce additional noise that interferes with model judgment. This finding offers important insights for the optimized design of task descriptions in subsequent work, particularly when handling generative tasks, where the applicability and formulation of task descriptions require careful reconsideration.
RQ5: How does data scale affect model performance?
In this section, we investigate the impact of fine-tuning data scale on model performance. Using stratified sampling, we created three subsets of training data at 25%, 50%, and 75% proportions to fine-tune the GLM-4-9B. The model’s performance was then evaluated on the test set, with experimental results shown in Figure 4(b). Our findings reveal that:
$\bullet$ The model’s performance across different tasks exhibits a positive correlation with the amount of fine-tuning data, though the sensitivity to data volume varies significantly between tasks. Specifically, the pUII metric demonstrates steady improvement as data scale increases, whereas the performance of the UIU task remains largely insensitive to variations in data quantity. This finding highlights the differential data requirements imposed by task characteristics, suggesting the need for further research on optimizing the composition and scale of instruction data based on task-specific properties to achieve more efficient model training outcomes.
RQ6: What are the performance differences between fine-tuned models and closed-source LLMs on recommendation tasks?
In this section, we compared the performance differences between fine-tuned models and larger models. We selected the currently mainstream closed-source LLMs, GPT-3.5-Turbo (openai2022chatgpt) and Gemini 2.5 Flash (Gemini2025Comanici), as benchmark models and evaluated their performance on the UII and UIU tasks. The experimental results shown in Figure 5 demonstrate that:
$\bullet$ The fine-tuned models demonstrate superior performance on both UII and UIU tasks compared to the baseline model with larger parameter sizes. Notably, benefiting from the advantage of massive parameter scale, larger models still exhibits significantly better performance on the UII task than untuned open-source models. Further analysis reveals distinct task preference characteristics among different models: GPT-3.5-Turbo performs more effectively on UIU tasks, while Gemini 2.5 Flash demonstrates stronger processing capabilities on UII tasks, indicating fundamental differences in how different model architectures understand and solve tasks. This series of findings reaffirms the significant value of our developed ITDR dataset in recommendation tasks.
5. Conclusion
In this paper, we propose ITDR, an instruction tuning dataset constructed from two dimensions—user-item interaction and user-item understanding—effectively filling the gap of natural language datasets in recommendation systems. We systematically investigate leveraging ITDR to enhance LLMs’ performance on recommendation tasks. Through extensive experiments, we explore several key aspects, including task relationships, task description design, and the impact of data volume in instruction tuning, demonstrating ITDR’s effectiveness in enhancing LLM-based recommendations. We expect ITDR to advance research on LLMs for recommendation and stimulate further academic exploration of instruction tuning methodologies.
Limitations and Future Work
In this work, we construct a sufficient dataset for instruction tuning on recommendation systems and trained fine-tuned models based on this dataset. We identify the limitations of the current research and proposed directions for future improvements.
In terms of the dataset, although we integrat 7 types of recommendation subtasks from 13 source datasets, a large number of datasets and tasks remain unincorporated. Experimental results indicate that introducing more diverse data could further enhance the coverage and diversity of the dataset, which may positively impact model performance.
Regarding model parameters, our study only validat the effectiveness of models with parameters ranging from 3B to 9B. Future work could explore the impact of instruction tuning on larger-scale models (e.g., 13B, 70B-level) and conduct comparisons with existing closed-source LLMs, which would hold significant research value.
For the evaluation methods, due to resource constraints, we primarily focuse on average performance metrics at the task level. Subsequent research could analyze the interaction effects between different datasets at a finer granularity, providing more instructive recommendations for optimizing the composition of instruction tuning datasets.
Acknowledgements.
This work was supported in part by the National Key Research and Development Program of China under Grant (2023YFC3310700) and the National Natural Science Foundation of China (62202041).
Appendix A Details about Tasks and Datasets
A.1. User-Item Interaction
In recommendation systems, user-item interaction data serves as the core element for system optimization, capturing the multidimensional behavioral trajectories between users and recommended content. This includes explicit feedback (such as ratings and purchases) and implicit signals (such as clicks and browsing duration). These interaction data inherently reflect the complex mapping relationships between user preferences and item features, providing the foundational data for building accurate personalized recommendation models. Among these, how to effectively model user-item interaction relationships not only directly impacts the accuracy of recommendation results but also significantly influences user satisfaction. To this end, we design a series of subtasks focused on user-item interactions, aiming to enhance the capability of LLMs in capturing user behaviors and item features, with a particular emphasis on evaluating interaction modeling tasks across the following four tasks.
Rating Prediction (RP)
Rating prediction is a classic task in recommendation systems, which aims to predict a user’s potential rating for unrated items by analyzing their historical rating behaviors. We select four representative public datasets: MovieLens 32M (MovieLens), BookCrossing (Bookcrossing), Anime Dataset 2023111https://www.kaggle.com/datasets/dbdmobile/myanimelist-dataset, and Yelp (Yelp).
Top-K Recommendation (TKR)
Top-K recommendation aims to predict the top K items that a user is most likely to be interested in next by analyzing their historical interaction sequences. Unlike traditional rating prediction or click-through rate prediction, Top-K recommendation places greater emphasis on the ranking quality and diversity of the recommendation list. In this study, we utilize the following datasets:Amazon Review 2023 (AmazonReviews2023), MovieLens 32M (MovieLens), MicroLens (MicroLens), PixelRec (PixelRec), Amazon Books Reviews [^2] https://www.kaggle.com/datasets/mohamedbakhet/amazon-books-reviews, MIND (MIND), Anime Dataset 20231.
Cross-Domain Recommendation (CDR)
Cross-domain recommendation aims to enhance recommendation performance in a target domain by leveraging user behavior data from one or multiple source domains. Unlike single-domain recommendation, the core challenge of CDR lies in transferring user preferences and knowledge across different domains, particularly when facing data sparsity in the target domain. This study builds instruction data based on the Amazon Review 2023 (AmazonReviews2023) dataset, with the specific source and target domain division scheme detailed in Table 3.
Table 3. Selection strategies for source and target domains in cross-domain recommendation task.
Source Target
Beauty Fashion
CDs & Vinyl Video Games
Appliances Industrial & Scientific
Home & Kitchen Grocery & Gourmet Food
Next Item Recommendation (NIR)
Next item recommendation is a core task in sequential recommendation systems, with the primary objective of predicting the items a user is most likely to interact with next by modeling their historical interaction sequences (such as browsing, clicking, purchasing, etc.). We select the following publicly available datasets as data sources: Amazon Reviews 2023 (AmazonReviews2023), MicroLens (MicroLens), MovieLens 32M (MovieLens), and PixelRec (PixelRec).
A.2. User-Item Understanding
In recommendation systems, user-item understanding serves as the core dimension of system cognition. By deeply mining the bidirectional semantic associations between user characteristics and item attributes, it provides critical support for personalized recommendations. This process involves in-depth analysis at two levels: On the user side, the system constructs precise user profiles by analyzing historical behavioral data; on the item side, it focuses on analyzing target audience characteristics to form comprehensive item profiles. These multidimensional understandings establish mapping relationships between users and items, not only improving recommendation accuracy but also laying the cognitive foundation for explainable recommendation logic. Based on this, we design three user-item interaction subtasks aimed at enhancing the large language model’s joint reasoning and matching capabilities for user attributes and item features.
User Attribute Prediction (UAP)
User attribute prediction is a crucial task in building user profiles for recommendation systems. Its primary objective is to infer user’s static demographic characteristics or social attribute information by analyzing their interaction behaviors and preference patterns with items. Unlike traditional recommendation tasks, user attribute prediction does not aim to predict items that users might prefer, but rather deduces users’ inherent attribute features from their behavioral data. For this subtask, we select the following datasets as data sources: Last.FM 1/360K (Last.FM1K_360K), Anime Dataset 20231, BookCrossing (Bookcrossing), and MovieLens 1M (MovieLens).
Interest Recognition (IR)
The primary objective of interest recognition is to analyze user‘s historical interaction data to identify and summarize their preferences, thereby enabling more accurate personalized services. For this subtask, we select the following datasets as data sources: Amazon Books Reviews2, Anime Dataset 20231, and Steam (steam).
Target User Identification (TUI)
The core objective of target user identification is to analyze the attribute characteristics of items in order to identify the most likely potential user groups interested in those items. Unlike the traditional ”user-to-item” recommendation logic, target user identification achieves a reverse ”item-to-user” recommendation mechanism. For this subtask, we select the following datasets as data sources: Amazon Reviews 2023 (AmazonReviews2023), BookCrossing (Bookcrossing), Anime Dataset 20231, Steam (steam), Yelp (Yelp), MicroLens (MicroLens), PixelRec (PixelRec), MovieLens 32M (MovieLens), and MIND (MIND).
Appendix B Statistics of ITDR and Additional Results
Detailed statistics of ITDR and experimental results are presented in Table 5 – Table 17.
Table 5 summarizes the complete statistical information of ITDR, encompassing the following dimensions: subtasks, data sources, evaluation metrics, sample size, and the average input/output tokens calculated based on the DeepSeek-v3 tokenizer.
According to the experimental data in Table 6 – Table 8, the fine-tuned model shows significant improvements in performance metrics across all tasks, particularly in sequence recommendation tasks (TKR, CDR, NIR) and generation tasks (IR, TUI). The results fully validate the effectiveness of ITDR in improving the performance of LLMs in recommendation tasks.
Based on the experimental results of the ablation studies at the root tasks shown in Table 9 – Table 11, we can draw the following conclusions: first, the UIU tasks enhance the performance of the UII tasks, which aligns with our expectations. This indicates that the knowledge learned by the LLMs through the UIU tasks can be effectively transferred and guide the reasoning process of the UII tasks. However, the study also revealed a noteworthy phenomenon: when the model is fine-tuned using only the UII data, the performance of the UIU tasks exhibits a clear decline. This counterintuitive observation suggests that there may be complex interaction mechanisms between different task datasets, requiring further research to explore their underlying relationships. At the same time, this finding provides important empirical evidence and research directions for optimizing data composition strategies in future model fine-tuning processes.
Table 12 – Table 14 present the experimental results of GLM-4-9B in subtask level ablation studies. We observe interaction effects among different subtasks. Specifically, when the CDR task is removed during fine-tuning, the model’s performance on this task surpassed that of the unfine-tuned baseline results (as shown in Table 6). This indicates that the knowledge learned from other subtasks can be effectively transferred to this task, positively enhancing its reasoning capabilities. However, such interaction effects are not always beneficial. For instance, when the RP task is removed during fine-tuning, the model’s average performance on this task is lower than the baseline level. These findings highlight the importance of formulating rational data selection strategies during the model fine-tuning process.
Based on the experimental results from Table 15 – Table 17, GPT-3.5-Turbo and Gemini 2.5 Flash exhibit the following characteristics in UII and UIU tasks: First, owing to their substantial model parameters, these two LLMs demonstrate significantly superior performance in UII tasks compared to untuned small-scale models, though they still fall short of the level achieved by small models that have undergone instruction tuning. Second, in UIU tasks, their performance shows no significant difference when compared to untuned small-scale models. Notably, the two models exhibit performance divergence across different tasks: GPT-3.5-Turbo performs more prominently in UIU tasks, while Gemini 2.5 Flash shows a relative advantage in UII tasks. This phenomenon suggests that LLMs with different architectures may possess distinct strengths in specific task domains.
Appendix C Data Examples
Data examples for each subtask are presented in Table 18 – Table 24. The construction process for each instruction is shown below.
Taking the instruction data from the Anime Dataset 2023 shown in Table 18 as an example, we adopt the following method to construct instruction data for the RP task: First, valid user samples with no fewer than 11 rating records are filtered. Then, the viewing records of each user are sorted in ascending chronological order. Next, the first 10 rating records are extracted as the model’s prior knowledge input. Finally, the 11th rating record is used as the prediction target, with its true rating value serving as the ground truth for model performance evaluation.
Table 19 presents the TKR task instruction construction method based on the PixelRec data source. We first filter user samples with no fewer than 11 interaction records and sort their interaction sequences in chronological order. When constructing instructions, the first 5 interaction records of each user are used as prior knowledge input for the model. Meanwhile, the subsequent 5 positive samples are mixed with 5 randomly sampled negative samples and shuffled to form a candidate set to be ranked. The output content is the correct ranking order of the samples to be sorted.
When constructing instructions for the CDR task based on the Amazon Review 2023 dataset, we implement the following process: first, we extract the user’s first 10 interaction records sorted in ascending chronological order from the source domain as prior knowledge input for the model; second, we select the top 5 positive sample records from the target domain, sorted in ascending order by time, and combine them with 5 randomly sampled negative records to form the candidate ranking set. The model’s expected output is the correct sorted order of these samples. Table 20 shows the details about the instruction.
Table 21 presents an instruction example for the NIR task, which follows a design approach similar to the TKR task. We first filter user samples with no fewer than 6 interaction records and sort their interaction sequences in chronological order. The first 5 interaction records of each user serve as prior knowledge input for the model, while the remaining one positive sample is mixed with 4 randomly sampled negative samples and shuffled to form the candidate set for prediction. The model’s objective is to accurately identify the positive sample from the candidate set.
Taking the MovieLens 1M dataset as an example, we designed an instruction data construction method for the UAP task: First, we filtered out each user’s movie-watching records with ratings above 3 as positive samples and sorted them in ascending order by viewing time. Then, we used the user’s static attribute features as the target variables for the model to predict. As shown in Table 22, we provide specific instruction examples to illustrate the data construction process.
Table 23 presents an example of instruction construction for IR tasks, which is built based on the Amazon Books Reviews dataset. We first categorize user ratings into ”like” (greater than 3 points) and ”dislike” (less than 3 points) using a threshold of 3, then select 3 representative samples from each category, including book titles and their corresponding reviews, as prior knowledge for the model. Since this is a generative task and the original dataset lacks prediction targets, we employ the DeepSeek-V3 to generate reliable ground truth data as an evaluation benchmark, thereby providing high-quality data support for model training and evaluation.
Table 24 demonstrates an example of constructing TUI task instructions based on the Steam dataset. This construction process includes the following steps: First, for each game, extract its key attributes, including feature information such as the game’s name, genre, tags, and gameplay. Second, randomly sample real player reviews of the game (which may contain negative feedback). Similar to IR tasks, we use DeepSeek-V3 to generate standard answer data as the expected output for the instructions.
Appendix D Notations and Terminologies
We summarize the key notations used in this paper in Table 4.
Table 4. Key notations.
Symbol Description
UII User-Item Interaction
UIU User-Item Understanding
RP Rating Prediction
TKR Top-K Recommendation
CDR Cross-Domain Recommendation
NIR Next Item Recommendation
UAP User Attribute Prediction
IR Interest Recognition
TUI Target User Identification
$ST$ set of subtasks
$M$ set of evaluation metrics for sequential recommendation tasks
$D_{st}$ set of all subtasks data
$pUII$ performance of user-item interaction
$pUIU$ performance of user-item understanding
$D_{UAP}$ set of UAP task
$M^{\prime}$ set of evaluation metrics for IR and TUI tasks
Table 5. The statistics of all datasets.
Task Dataset Metrics Examples Avg In Avg Out
Rating Prediction MovieLens 32M RMSE 6,250 162.32 3.00
Rating Prediction BookCrossing 4,623 78.47 1.00
Rating Prediction Anime Dataset 2023 6,250 120.02 1.00
Rating Prediction Yelp 5,707 323.98 1.00
Top-K Recommendation Amazon Reviews 2023
N@10
H@1
P@5
5,276 435.32 264.53
Top-K Recommendation MovieLens 32M 6,250 208.35 109.13
Top-K Recommendation MicroLens 6,250 325.37 186.57
Top-K Recommendation PixelRec 6,250 357.41 206.69
Top-K Recommendation Amazon Books Reviews 6,250 216.37 118.45
Top-K Recommendation MIND 6,250 314.42 177.28
Top-K Recommendation Anime Dataset 2023 6,250 194.19 105.22
Cross-Domain Recommendation Amazon Reviews 2023
N@10
H@1
P@5
4,118 651.50 313.67
Next Item Recommendation Amazon Reviews 2023 H@1 5,830 291.35 24.10
Next Item Recommendation MicroLens 6,250 251.32 14.70
Next Item Recommendation MovieLens 32M 6,250 306.52 9.34
Next Item Recommendation PixelRec 6,250 251.48 18.30
User Attribute Prediction Last.FM 1K
ACC(country)
ACC(gender)
RMSE(age)
943 590.56 11.44
User Attribute Prediction Last.FM 360K
ACC(country)
ACC(gender)
RMSE(age)
6,250 267.11 12.72
User Attribute Prediction Anime Dataset 2023
ACC(location)
ACC(gender)
RMSE(Birth year)
6,250 571.91 16.55
User Attribute Prediction BookCrossing
ACC(location)
RMSE(age)
6,250 71.81 14.30
User Attribute Prediction MovieLens 1M
ACC(age group)
ACC(gender)
ACC(occupation)
6,040 310.24 18.30
Interest Recognition Amazon Books Reviews
BLEU-1&2,
ROUGE-1&2&L
6,250 1,675.46 107.18
Interest Recognition Anime Dataset 2023 6,250 565.25 54.96
Interest Recognition Steam 6,228 802.22 358.40
Target User Identification Amazon Reviews 2023
BLEU-1&2
ROUGE-1&2&L
6,300 370.89 253.21
Target User Identification BookCrossing 6,250 86.20 226.82
Target User Identification Anime Dataset 2023 6,250 157.67 218.72
Target User Identification Steam 6,250 475.94 250.80
Target User Identification Yelp 6,250 573.26 249.21
Target User Identification MicroLens 6,250 53.73 218.82
Target User Identification PixelRec 6,250 97.30 230.21
Target User Identification MovieLens 32M 6,250 122.81 238.53
Target User Identification MIND 6,250 95.10 166.82
  • *The average input/output tokens are calculated using the DeepSeek-v3 tokenizer.
Table 6. Performance on the user-item interaction tasks.
GLM-4 Qwen2.5-Ins Qwen2.5 LLaMA-3.2
Task & Dataset Metric No FT FT 25% 50% 75% No FT FT No FT FT No FT FT
Rating Prediction
MovieLens 32M RMSE 1.1469 1.0339 0.9945 1.0780 1.0010 1.1151 1.0276 1.0108 1.0877 1.1239 1.1050
BookCrossing RMSE 1.8562 1.7430 1.7289 1.7477 1.7242 1.9004 1.8518 1.9546 1.7754 1.9151 2.0589
Anime Dataset 2023 RMSE 1.5969 1.4213 1.5179 1.4085 1.3813 1.5356 1.6050 1.5526 1.5466 1.6952 1.5543
Yelp RMSE 2.1065 2.0196 1.9206 1.9779 2.0022 2.3952 2.0562 1.9217 2.0011 2.2891 2.4324
Top-K Recommendation
Amazon Reviews 2023 N@10 0.6976 0.9191 0.9011 0.9099 0.9079 0.4778 0.9114 0.7201 0.9151 0.4298 0.8724
H@1 0.4958 0.8644 0.8305 0.8432 0.8220 0.4831 0.8559 0.6441 0.8517 0.3347 0.8136
P@5 0.4573 0.7805 0.7466 0.7644 0.7669 0.3525 0.7508 0.5263 0.7695 0.2973 0.7364
MovieLens 32M N@10 0.6831 0.9890 0.9891 0.9868 0.9900 0.7100 0.9889 0.7977 0.9914 0.4949 0.9911
H@1 0.5640 0.9880 0.9880 0.9800 0.9920 0.7040 0.9800 0.7080 0.9960 0.4320 0.9960
P@5 0.4616 0.9608 0.9608 0.9600 0.9632 0.5832 0.9640 0.6512 0.9640 0.3728 0.9648
MicroLens N@10 0.6784 0.8557 0.8507 0.8463 0.8620 0.5198 0.8491 0.6861 0.8620 0.4208 0.8216
H@1 0.5520 0.6960 0.7160 0.7000 0.7440 0.4400 0.7080 0.6120 0.7680 0.3640 0.7000
P@5 0.4434 0.6600 0.6384 0.6400 0.6560 0.3448 0.6424 0.4600 0.6416 0.2947 0.6416
PixelRec N@10 0.6718 0.8386 0.8287 0.8287 0.8320 0.5617 0.8244 0.5596 0.8435 0.3749 0.7999
H@1 0.5160 0.6640 0.6560 0.6600 0.6480 0.4640 0.6520 0.4200 0.7080 0.3320 0.6960
P@5 0.4432 0.6256 0.6016 0.6072 0.6136 0.3544 0.5936 0.3416 0.6216 0.2576 0.6096
Amazon Books Reviews N@10 0.8796 0.9841 0.9857 0.9828 0.9851 0.4745 0.9832 0.8133 0.9819 0.5640 0.9828
H@1 0.8360 0.9840 0.9840 0.9680 0.9920 0.5680 0.9800 0.8560 0.9680 0.5760 0.9920
P@5 0.6712 0.9400 0.9344 0.9352 0.9352 0.3800 0.9368 0.6520 0.9376 0.4365 0.9336
MIND N@10 0.6976 0.8355 0.8197 0.8392 0.8414 0.6073 0.8225 0.4693 0.8337 0.3510 0.7926
H@1 0.4960 0.6600 0.6120 0.6720 0.6920 0.5440 0.6120 0.3240 0.6720 0.2720 0.6480
P@5 0.4574 0.6136 0.5968 0.6136 0.6256 0.3760 0.6072 0.3925 0.6000 0.2162 0.5912
Anime Dataset 2023 N@10 0.7284 0.9743 0.9624 0.9673 0.9732 0.7679 0.9662 0.7108 0.9688 0.4871 0.9771
H@1 0.6840 0.9600 0.9400 0.9560 0.9520 0.8040 0.9720 0.6240 0.9600 0.4560 0.9840
P@5 0.4856 0.9240 0.8968 0.9056 0.9192 0.5808 0.8808 0.4944 0.9016 0.3525 0.9280
Cross-Domain Recommendation
Amazon Reviews 2023 N@10 0.5983 0.9086 0.8888 0.8931 0.9012 0.3855 0.8781 0.5376 0.8822 0.4339 0.8932
H@1 0.5449 0.8503 0.8024 0.8263 0.8443 0.3353 0.7784 0.4012 0.7784 0.3832 0.8683
P@5 0.4968 0.7449 0.7054 0.7126 0.7222 0.2599 0.7030 0.3401 0.7066 0.3110 0.7377
Next Item Recommendation
Amazon Reviews 2023 H@1 0.3247 0.7879 0.6883 0.7662 0.7835 0.4589 0.6407 0.3810 0.6537 0.2987 0.5108
MicroLens H@1 0.3320 0.9200 0.7320 0.8560 0.9040 0.3720 0.5920 0.3320 0.6320 0.2520 0.4040
MovieLens 32M H@1 0.3560 0.6080 0.5920 0.6120 0.5880 0.2240 0.5320 0.2720 0.5240 0.2360 0.4840
PixelRec H@1 0.1720 0.8760 0.7520 0.8280 0.8400 0.2360 0.5200 0.1800 0.6360 0.1520 0.3960
  • *“FT” means “fine-tuning”. “25%” indicates that the model was finetuned using 25% of the training data.
Table 7. Performance on the user-item understanding task.
GLM-4 Qwen2.5-Ins Qwen2.5 LLaMA-3.2
Task & Dataset Metric No FT FT 25% 50% 75% No FT FT No FT FT No FT FT
User Attribute Prediction
Last.FM 1K ACC(c) 0.1892 0.4865 0.5676 0.5135 0.5405 0.2162 0.4865 0.1622 0.5676 0.1622 0.3514
ACC(g) 0.2702 0.5405 0.5946 0.6216 0.5946 0.6216 0.6757 0.2432 0.4054 0.1081 0.2432
RMSE(a) 6.3650 6.1489 6.5826 6.3622 6.2469 5.9780 5.7680 6.3690 6.1056 5.9512 6.0075
Last.FM 360K ACC(c) 0.2560 0.4720 0.4560 0.4400 0.4480 0.2440 0.3840 0.1480 0.3880 0.0760 0.4120
ACC(g) 0.5480 0.7240 0.7280 0.7440 0.7200 0.7080 0.7320 0.5320 0.6960 0.5800 0.7160
RMSE(age) 10.2369 8.5492 8.8794 9.0100 9.0044 9.1571 9.2065 9.1696 9.1297 9.4161 8.9618
Anime Dataset 2023 ACC(location) 0.0000 0.0440 0.0240 0.0600 0.0600 0.0040 0.0320 0.0040 0.0120 0.0080 0.0280
ACC(gender) 0.6400 0.8120 0.8000 0.8320 0.8160 0.5240 0.6600 0.5400 0.6840 0.5640 0.7160
RMSE(Birth year) 5.9435 4.5677 4.6994 4.4708 4.7019 6.2404 5.7372 6.8963 4.8662 6.1399 5.0570
BookCrossing ACC(location) 0.0000 0.0520 0.0760 0.0520 0.0720 0.0280 0.0560 0.0120 0.0560 0.0280 0.0520
RMSE(age) 18.9861 15.9227 15.8465 16.0752 16.1434 15.0502 15.2591 15.6593 15.9636 16.0585 15.7151
MovieLens 1M ACC(age group) 0.3458 0.3958 0.3958 0.3833 0.3583 0.3167 0.3250 0.3167 0.3708 0.3125 0.3625
ACC(gender) 0.5292 0.7708 0.7375 0.7792 0.7542 0.4167 0.7333 0.4958 0.7375 0.4792 0.7083
ACC(occupation) 0.0000 0.1375 0.1250 0.1417 0.1083 0.0042 0.0958 0.0083 0.1042 0.0042 0.1125
Interest Recognition
Amazon Books Reviews BLEU-1 0.1436 0.3514 0.3377 0.3539 0.3471 0.0807 0.3268 0.1200 0.3118 0.0636 0.3188
BLEU-2 0.0564 0.2161 0.2061 0.2208 0.2117 0.0362 0.1934 0.0488 0.1863 0.0243 0.1920
ROUGE-1 0.1884 0.4138 0.4028 0.4145 0.4096 0.1698 0.3649 0.1789 0.3671 0.1065 0.3690
ROUGE-2 0.0314 0.1711 0.1663 0.1788 0.1691 0.0312 0.1352 0.0306 0.1371 0.0132 0.1401
ROUGE-L 0.1321 0.3878 0.3762 0.3873 0.3839 0.0842 0.2742 0.1075 0.2664 0.0336 0.2710
Anime Dataset 2023 BLEU-1 0.1500 0.4440 0.4166 0.4332 0.4361 0.0566 0.4335 0.0419 0.3422 0.0539 0.3916
BLEU-2 0.0722 0.3379 0.3052 0.3190 0.3220 0.0274 0.3074 0.0209 0.2398 0.0133 0.2751
ROUGE-1 0.2299 0.5396 0.5244 0.5356 0.5303 0.1322 0.5084 0.1423 0.4651 0.1037 0.4874
ROUGE-2 0.0562 0.3069 0.2748 0.2896 0.2860 0.0272 0.2543 0.0311 0.2242 0.0072 0.2412
ROUGE-L 0.1496 0.5052 0.4854 0.4980 0.4950 0.0729 0.4331 0.0542 0.3562 0.0343 0.3916
Steam BLEU-1 0.1526 0.4613 0.4467 0.4623 0.4602 0.3041 0.4450 0.1830 0.4056 0.0736 0.3587
BLEU-2 0.0796 0.3109 0.2921 0.3105 0.3113 0.1590 0.2932 0.0966 0.2737 0.0232 0.2363
ROUGE-1 0.2392 0.4938 0.4765 0.4919 0.4951 0.3349 0.4637 0.2788 0.4519 0.1504 0.4288
ROUGE-2 0.0691 0.2338 0.2152 0.2323 0.2352 0.0978 0.2098 0.0827 0.2210 0.0311 0.1970
ROUGE-L 0.1539 0.4616 0.4428 0.4592 0.4618 0.1930 0.3003 0.1544 0.3023 0.0369 0.2760
  • *“FT” means “fine-tuning”. “25%” indicates that the model was finetuned using 25% of the training data.
Table 8. Continued from Table 7.
GLM-4 Qwen2.5-Ins Qwen2.5 LLaMA-3.2
Task & Dataset Metric No FT FT 25% 50% 75% No FT FT No FT FT No FT FT
Target User Identification
Amazon Reviews 2023 BLEU-1 0.2039 0.4819 0.4782 0.4856 0.4872 0.3008 0.4751 0.2082 0.4566 0.1254 0.4284
BLEU-2 0.1088 0.3314 0.3252 0.3338 0.3341 0.1730 0.3240 0.1149 0.3122 0.0606 0.2892
ROUGE-1 0.3118 0.5152 0.5059 0.5130 0.5151 0.3730 0.4875 0.3157 0.4838 0.2110 0.4782
ROUGE-2 0.0982 0.2552 0.2463 0.2537 0.2549 0.1262 0.2393 0.1036 0.2399 0.0630 0.2330
ROUGE-L 0.1912 0.4876 0.4770 0.4843 0.4862 0.2009 0.3246 0.1610 0.3139 0.0573 0.3020
BookCrossing BLEU-1 0.1960 0.5119 0.5035 0.5028 0.5075 0.3287 0.4905 0.1621 0.4751 0.0944 0.4120
BLEU-2 0.1093 0.3611 0.3518 0.3497 0.3573 0.1875 0.3342 0.0873 0.3319 0.0435 0.2870
ROUGE-1 0.3200 0.5422 0.5341 0.5306 0.5371 0.3841 0.5073 0.3024 0.5000 0.2138 0.4956
ROUGE-2 0.1074 0.2783 0.2687 0.2666 0.2750 0.1280 0.2455 0.0921 0.2557 0.0722 0.2536
ROUGE-L 0.1946 0.5099 0.5024 0.4993 0.5065 0.2129 0.3414 0.1424 0.3425 0.0513 0.3134
Anime Dataset 2023 BLEU-1 0.1356 0.5438 0.5372 0.5412 0.5513 0.3330 0.5322 0.1874 0.5177 0.0421 0.4171
BLEU-2 0.0749 0.4036 0.3901 0.3970 0.4087 0.1862 0.3867 0.1023 0.3818 0.0032 0.3069
ROUGE-1 0.2932 0.5835 0.5680 0.5759 0.5863 0.3791 0.5502 0.3243 0.5462 0.2123 0.5482
ROUGE-2 0.0948 0.3256 0.3052 0.3152 0.3272 0.1215 0.3021 0.1007 0.3072 0.0662 0.3057
ROUGE-L 0.1824 0.5513 0.5347 0.5425 0.5519 0.2128 0.3898 0.1618 0.3887 0.0334 0.3449
Steam BLEU-1 0.1854 0.5086 0.5046 0.5044 0.5115 0.2777 0.4996 0.2219 0.4562 0.0959 0.4297
BLEU-2 0.0984 0.3660 0.3563 0.3590 0.3657 0.1578 0.3536 0.1254 0.3276 0.0431 0.3068
ROUGE-1 0.3018 0.5381 0.5292 0.5316 0.5368 0.3526 0.5056 0.3281 0.4862 0.1958 0.4876
ROUGE-2 0.0917 0.2840 0.2687 0.2743 0.2811 0.1117 0.2645 0.1073 0.2621 0.0547 0.2571
ROUGE-L 0.1870 0.5081 0.4961 0.4994 0.5051 0.1928 0.3622 0.1828 0.3508 0.0530 0.3334
Yelp BLEU-1 0.2594 0.4829 0.4742 0.4832 0.4798 0.2673 0.4683 0.2291 0.4465 0.1370 0.4273
BLEU-2 0.1465 0.3297 0.3194 0.3284 0.3266 0.1502 0.3109 0.1311 0.3023 0.0737 0.2891
ROUGE-1 0.3486 0.5166 0.5084 0.5150 0.5157 0.3585 0.4797 0.3344 0.4865 0.2224 0.4810
ROUGE-2 0.1171 0.2547 0.2441 0.2513 0.2510 0.1140 0.2247 0.1149 0.2382 0.0734 0.2356
ROUGE-L 0.1972 0.4885 0.4812 0.4859 0.4858 0.1885 0.3186 0.1781 0.3231 0.0654 0.3136
MicroLens BLEU-1 0.1003 0.4998 0.4934 0.4966 0.5053 0.3549 0.4849 0.1410 0.4542 0.0248 0.3527
BLEU-2 0.0612 0.3583 0.3503 0.3556 0.3630 0.2133 0.3389 0.0851 0.3247 0.1059 0.2478
ROUGE-1 0.2884 0.5373 0.5293 0.5345 0.5393 0.3919 0.5048 0.3227 0.4983 0.2154 0.5021
ROUGE-2 0.1082 0.2826 0.2728 0.2812 0.2851 0.1459 0.2572 0.1196 0.2673 0.0774 0.2617
ROUGE-L 0.1727 0.5082 0.4999 0.5051 0.5108 0.2342 0.3487 0.1545 0.3526 0.0291 0.2966
PixelRec BLEU-1 0.2107 0.4879 0.4810 0.4839 0.4851 0.3412 0.4694 0.1810 0.4606 0.0819 0.3706
BLEU-2 0.1234 0.3472 0.3373 0.3413 0.3424 0.1915 0.3243 0.1009 0.3234 0.0344 0.2611
ROUGE-1 0.3359 0.5246 0.5161 0.5201 0.5224 0.3746 0.4941 0.3152 0.4951 0.2272 0.4950
ROUGE-2 0.1203 0.2696 0.2561 0.2625 0.2648 0.1236 0.2437 0.1024 0.2540 0.0795 0.2540
ROUGE-L 0.1983 0.4935 0.4832 0.4871 0.4890 0.2092 0.3320 0.1560 0.3385 0.0490 0.3012
MovieLens 32M BLEU-1 0.1527 0.5260 0.5151 0.5258 0.5209 0.3395 0.5194 0.1770 0.4813 0.0578 0.4026
BLEU-2 0.0876 0.3841 0.3697 0.3802 0.3808 0.1900 0.3698 0.0939 0.3476 0.0141 0.2887
ROUGE-1 0.2958 0.5513 0.5384 0.5473 0.5477 0.3709 0.5231 0.2836 0.5107 0.2111 0.5172
ROUGE-2 0.1022 0.2982 0.2818 0.2917 0.2974 0.1200 0.2764 0.0867 0.2773 0.0701 0.2800
ROUGE-L 0.1779 0.5126 0.4994 0.5081 0.5087 0.2036 0.3691 0.1435 0.3610 0.0354 0.3225
MIND BLEU-1 0.1190 0.5575 0.5401 0.5536 0.5507 0.3534 0.5323 0.1429 0.5357 0.0485 0.3386
BLEU-2 0.0723 0.4045 0.3833 0.3960 0.3955 0.2046 0.3738 0.0826 0.3851 0.0581 0.2374
ROUGE-1 0.3457 0.6106 0.5956 0.6064 0.6061 0.4210 0.5670 0.3591 0.5763 0.2124 0.5415
ROUGE-2 0.1259 0.3216 0.3015 0.3114 0.3131 0.1400 0.2840 0.1191 0.3040 0.0577 0.2723
ROUGE-L 0.2101 0.5788 0.5652 0.5746 0.5752 0.2325 0.3853 0.1409 0.4008 0.0123 0.2951
Table 9. Performance on user-item interaction tasks after removing different root tasks using ITDR across four models.
GLM-4 Qwen2.5-Ins Qwen2.5 LLaMA-3.2
Task & Dataset Metric w/o UII w/o UIU w/o UII w/o UIU w/o UII w/o UIU w/o UII w/o UIU
Rating Prediction
MovieLens 32M RMSE 0.9381 1.0856 1.1694 1.0761 1.0228 1.0954 1.1422 1.2661
BookCrossing RMSE 1.7941 1.7846 2.3248 1.9491 1.9047 1.8163 1.7800 1.7678
Anime Dataset 2023 RMSE 1.5681 1.3856 1.7760 1.5569 1.6997 1.4642 1.6383 1.4339
Yelp RMSE 1.9365 1.9768 2.5018 2.0077 2.0514 1.9978 2.1714 2.0477
Top-K Recommendation
Amazon Reviews 2023 N@10 0.7248 0.9128 0.7323 0.9121 0.7508 0.8971 0.4411 0.8103
H@1 0.5254 0.8262 0.6186 0.8432 0.7161 0.8136 0.3474 0.7796
P@5 0.4894 0.7703 0.5153 0.7653 0.5412 0.7523 0.3115 0.6961
MovieLens 32M N@10 0.7217 0.9841 0.8562 0.9899 0.8152 0.9900 0.5173 0.9902
H@1 0.6080 0.9720 0.8520 0.9840 0.7760 0.9880 0.4920 0.9920
P@5 0.4864 0.9536 0.6696 0.9616 0.6815 0.9550 0.3987 0.9123
MicroLens N@10 0.7024 0.8433 0.6758 0.8635 0.7132 0.8023 0.4399 0.7714
H@1 0.5800 0.6360 0.5600 0.7720 0.6560 0.7400 0.4320 0.6040
P@5 0.4720 0.6496 0.4744 0.6632 0.4833 0.5869 0.3264 0.6103
PixelRec N@10 0.7044 0.8019 0.6488 0.8374 0.6273 0.8277 0.4386 0.7402
H@1 0.5400 0.5640 0.4760 0.6880 0.5320 0.6440 0.3560 0.6760
P@5 0.5026 0.6072 0.4408 0.6168 0.4184 0.6064 0.3280 0.5928
Amazon Books Reviews N@10 0.9053 0.9862 0.7568 0.9826 0.8523 0.9755 0.6089 0.9266
H@1 0.8600 0.9640 0.7400 0.9760 0.8800 0.9520 0.6120 0.9760
P@5 0.7452 0.9400 0.5496 0.9360 0.7059 0.9011 0.4392 0.9352
MIND N@10 0.7311 0.8348 0.6841 0.8382 0.7103 0.8219 0.4094 0.7328
H@1 0.5200 0.6560 0.4480 0.6680 0.4120 0.6680 0.3000 0.5960
P@5 0.4926 0.6184 0.4568 0.6272 0.4584 0.5741 0.2927 0.5513
Anime Dataset 2023 N@10 0.7651 0.9732 0.7697 0.9786 0.7737 0.9780 0.5017 0.9327
H@1 0.7360 0.9440 0.7400 0.9760 0.7400 0.9240 0.4800 0.9680
P@5 0.5736 0.9272 0.5384 0.9232 0.5976 0.9200 0.4351 0.9248
Cross-Domain Recommendation
Amazon Reviews 2023 N@10 0.6632 0.9021 0.6835 0.8943 0.5831 0.8544 0.5585 0.8493
H@1 0.6107 0.8383 0.5329 0.7844 0.4311 0.6886 0.4431 0.8443
P@5 0.5819 0.7329 0.4817 0.7389 0.3914 0.6819 0.3976 0.6825
Next Item Recommendation
Amazon Reviews 2023 H@1 0.3723 0.7532 0.3203 0.6840 0.4112 0.6017 0.3160 0.4762
MicroLens H@1 0.4000 0.8760 0.2840 0.6120 0.3840 0.5800 0.2840 0.3800
MovieLens 32M H@1 0.4120 0.5640 0.3080 0.6240 0.3080 0.4960 0.3200 0.4600
PixelRec H@1 0.2600 0.8200 0.2080 0.7360 0.1920 0.6000 0.1600 0.2400
  • *“UII” means “user-item interaction”, “UIU” means “user-item understanding”.
Table 10. Performance on user-item understanding tasks after removing different root tasks using ITDR across four models.
GLM-4 Qwen2.5-Ins Qwen2.5 LLaMA-3.2
Task & Dataset Metric w/o UII w/o UIU w/o UII w/o UIU w/o UII w/o UIU w/o UII w/o UIU
User Attribute Prediction
Last.FM 1K ACC(country) 0.5946 0.2432 0.4595 0.2973 0.5405 0.2162 0.3243 0.1351
ACC(gender) 0.5135 0.1081 0.7027 0.6486 0.3784 0.1892 0.2162 0.1081
RMSE(age) 6.2560 6.8260 6.0467 6.4431 7.3695 7.6896 6.5574 6.9862
Last.FM 360K ACC(country) 0.4440 0.1040 0.3880 0.2640 0.3920 0.2600 0.4080 0.0960
ACC(gender) 0.7440 0.2720 0.7160 0.7240 0.7000 0.4080 0.7200 0.4680
RMSE(age) 8.9492 9.5057 9.1571 8.6799 8.7848 8.7544 8.9637 8.9800
Anime Dataset 2023 ACC(location) 0.0480 0.0080 0.0520 0.0080 0.0320 0.0080 0.0440 0.0080
ACC(gender) 0.8160 0.6840 0.6560 0.6520 0.6520 0.5600 0.7080 0.6320
RMSE(Birth year) 4.4940 5.8546 6.0712 9.0540 5.3100 5.6477 5.2917 6.8521
BookCrossing ACC(location) 0.0520 0.0000 0.0680 0.0040 0.0640 0.0080 0.0600 0.0040
RMSE(age) 15.6484 18.0959 15.2495 18.1059 15.6144 17.8786 15.5041 18.0268
MovieLens 1M ACC(age group) 0.3833 0.1667 0.3333 0.2292 0.3750 0.2500 0.3583 0.2125
ACC(gender) 0.7292 0.7542 0.7250 0.7156 0.7125 0.7417 0.7208 0.7333
ACC(occupation) 0.1167 0.0000 0.1000 0.0083 0.0875 0.0042 0.1014 0.0000
Interest Recognition
Amazon Books Reviews BLEU-1 0.3577 0.0015 0.3333 0.1618 0.2692 0.0207 0.2204 0.0241
BLEU-2 0.2239 0.0005 0.2011 0.0690 0.1569 0.0082 0.1294 0.0084
ROUGE-1 0.4182 0.0293 0.3711 0.1971 0.3605 0.1711 0.3594 0.0595
ROUGE-2 0.1807 0.0037 0.1420 0.0393 0.1289 0.0274 0.1286 0.0068
ROUGE-L 0.3925 0.0283 0.2786 0.1433 0.2368 0.0297 0.2090 0.0361
Anime Dataset 2023 BLEU-1 0.4339 0.1530 0.4362 0.2029 0.2916 0.0208 0.2093 0.0943
BLEU-2 0.3291 0.0781 0.3167 0.0994 0.1960 0.0090 0.1414 0.0475
ROUGE-1 0.5442 0.2918 0.5189 0.3066 0.4298 0.1238 0.4424 0.1927
ROUGE-2 0.3063 0.0764 0.2733 0.0760 0.1905 0.0212 0.1989 0.0479
ROUGE-L 0.5092 0.2488 0.4491 0.2173 0.3124 0.0305 0.2469 0.1057
Steam BLEU-1 0.4620 0.0000 0.4438 0.0240 0.3899 0.0403 0.3421 0.0260
BLEU-2 0.3128 0.0000 0.2930 0.0106 0.2601 0.0113 0.2253 0.0062
ROUGE-1 0.4958 0.0091 0.4612 0.1385 0.4486 0.0816 0.4335 0.0979
ROUGE-2 0.2366 0.0005 0.2115 0.0268 0.2122 0.0093 0.2018 0.0044
ROUGE-L 0.4634 0.0091 0.3024 0.0888 0.2904 0.0333 0.2635 0.0507
  • *“UII” means “user-item interaction”, “UIU” means “user-item understanding”.
Table 11. Continued from Table 10.
GLM-4 Qwen2.5-Ins Qwen2.5 LLaMA-3.2
Task & Dataset Metric w/o UII w/o UIU w/o UII w/o UIU w/o UII w/o UIU w/o UII w/o UIU
Target User Identification
Amazon Reviews 2023 BLEU-1 0.4884 0.0000 0.4707 0.0048 0.4359 0.0481 0.3675 0.0580
BLEU-2 0.3362 0.0000 0.3185 0.0023 0.2941 0.0196 0.2467 0.0178
ROUGE-1 0.5161 0.0395 0.4849 0.1146 0.4733 0.1410 0.4757 0.1345
ROUGE-2 0.2555 0.0066 0.2347 0.0283 0.2284 0.0285 0.2293 0.0169
ROUGE-L 0.4868 0.0373 0.3197 0.0692 0.3013 0.0484 0.2760 0.0905
BookCrossing BLEU-1 0.5110 0.0000 0.4980 0.0025 0.4637 0.0222 0.4130 0.0600
BLEU-2 0.3607 0.0000 0.3424 0.0009 0.3196 0.0065 0.2863 0.0191
ROUGE-1 0.5413 0.0338 0.5133 0.0654 0.4953 0.0743 0.4918 0.1539
ROUGE-2 0.2785 0.0076 0.2528 0.0124 0.2457 0.0074 0.2455 0.0183
ROUGE-L 0.5094 0.0326 0.3469 0.0433 0.3303 0.0252 0.3073 0.0927
Anime Dataset 2023 BLEU-1 0.5462 0.0000 0.5333 0.0020 0.5085 0.0100 0.4253 0.0654
BLEU-2 0.4045 0.0000 0.3863 0.0007 0.3669 0.0026 0.3113 0.0242
ROUGE-1 0.5839 0.0245 0.5514 0.0791 0.5394 0.0349 0.5426 0.1721
ROUGE-2 0.3251 0.0023 0.3005 0.0150 0.2921 0.0027 0.2990 0.0279
ROUGE-L 0.5506 0.0237 0.3920 0.0546 0.3750 0.0116 0.3428 0.1035
Steam BLEU-1 0.5135 0.0000 0.5037 0.0024 0.4329 0.0369 0.3790 0.0441
BLEU-2 0.3699 0.0000 0.3559 0.0011 0.3064 0.0137 0.2668 0.0147
ROUGE-1 0.5395 0.0379 0.5095 0.1041 0.4752 0.1303 0.4806 0.1474
ROUGE-2 0.2843 0.0083 0.2660 0.0200 0.2473 0.0196 0.2475 0.0218
ROUGE-L 0.5084 0.0362 0.3634 0.0612 0.3308 0.0398 0.3010 0.1004
Yelp BLEU-1 0.4815 0.0000 0.4697 0.0026 0.4368 0.0230 0.3586 0.0514
BLEU-2 0.3287 0.0000 0.3147 0.0008 0.2938 0.0066 0.2391 0.0133
ROUGE-1 0.5161 0.0385 0.4843 0.0799 0.4791 0.0838 0.4712 0.1373
ROUGE-2 0.2523 0.0061 0.2295 0.0120 0.2309 0.0074 0.2248 0.0128
ROUGE-L 0.4868 0.0372 0.3218 0.0501 0.3119 0.0267 0.2768 0.0857
MicroLens BLEU-1 0.5015 0.0004 0.4863 0.0002 0.4396 0.0180 0.3918 0.0234
BLEU-2 0.3611 0.0002 0.3412 0.0001 0.3115 0.0040 0.2734 0.0081
ROUGE-1 0.5366 0.0374 0.5074 0.0478 0.4972 0.0505 0.4874 0.1255
ROUGE-2 0.2840 0.0053 0.2603 0.0100 0.2590 0.0028 0.2466 0.0165
ROUGE-L 0.5066 0.0348 0.3520 0.0316 0.3395 0.0215 0.3075 0.0830
PixelRec BLEU-1 0.4942 0.0000 0.4646 0.0002 0.4525 0.0208 0.3885 0.0355
BLEU-2 0.3517 0.0000 0.3173 0.0001 0.3139 0.0060 0.2699 0.0115
ROUGE-1 0.5300 0.0450 0.4891 0.0402 0.4935 0.0700 0.4844 0.1270
ROUGE-2 0.2725 0.0061 0.2361 0.0077 0.2438 0.0075 0.2405 0.0136
ROUGE-L 0.4968 0.0410 0.3247 0.0284 0.3261 0.0247 0.3002 0.0855
MovieLens 32M BLEU-1 0.5270 0.0000 0.5126 0.0007 0.4737 0.0218 0.4467 0.0478
BLEU-2 0.3851 0.0000 0.3662 0.0003 0.3409 0.0049 0.3204 0.0108
ROUGE-1 0.5526 0.0221 0.5215 0.0513 0.5082 0.0723 0.5102 0.1447
ROUGE-2 0.2999 0.0018 0.2777 0.0043 0.2739 0.0042 0.2740 0.0103
ROUGE-L 0.5158 0.0217 0.3709 0.0336 0.3554 0.0235 0.3357 0.0979
MIND BLEU-1 0.5513 0.0000 0.5343 0.0004 0.5245 0.0250 0.3873 0.0770
BLEU-2 0.3959 0.0000 0.3754 0.0001 0.3701 0.0079 0.2696 0.0309
ROUGE-1 0.6057 0.0334 0.5681 0.0447 0.5699 0.0949 0.5681 0.2018
ROUGE-2 0.3132 0.0046 0.2845 0.0057 0.2877 0.0092 0.2868 0.0318
ROUGE-L 0.5760 0.0328 0.3846 0.0296 0.3803 0.0286 0.3172 0.1289
Table 12. Results for user-item interaction tasks on GLM-4-9B with ITDR removing different subtasks.
Task & Dataset Metric w/o RP w/o TKR w/o CDR w/o NIR w/o UAP w/o IR w/o TUI w/o TD
Rating Prediction
MovieLens 32M RMSE 1.0271 0.9975 1.0488 0.9022 1.0178 0.9854 0.9381 0.9592
BookCrossing RMSE 2.1443 1.6402 1.7131 1.7321 1.7020 1.7131 1.8445 1.7907
Anime Dataset 2023 RMSE 1.5336 1.3130 1.3900 1.3928 1.3638 1.3711 1.3784 1.4656
Yelp RMSE 2.2282 1.9331 1.9890 1.9702 2.0131 1.9354 2.0402 2.0251
Top-K Recommendation
Amazon Reviews 2023 N@10 0.9085 0.6419 0.9111 0.9037 0.9090 0.9156 0.9086 0.9078
H@1 0.8305 0.6017 0.8347 0.8051 0.8305 0.8602 0.8390 0.8136
P@5 0.7720 0.5034 0.7763 0.7551 0.7678 0.7797 0.7111 0.7712
MovieLens 32M N@10 0.9881 0.9439 0.9841 0.9903 0.9892 0.9849 0.9851 0.9892
H@1 0.9800 0.9560 0.9720 0.9840 0.9840 0.9760 0.9760 0.9880
P@5 0.9624 0.8360 0.9592 0.9680 0.9640 0.9576 0.9568 0.9592
MicroLens N@10 0.8567 0.6298 0.8548 0.8674 0.8582 0.8544 0.8516 0.8599
H@1 0.7080 0.5440 0.7080 0.7360 0.7160 0.7000 0.6880 0.7160
P@5 0.6608 0.4536 0.6536 0.6584 0.6680 0.6576 0.6448 0.6536
PixelRec N@10 0.8315 0.3460 0.8438 0.8391 0.8316 0.8274 0.8381 0.8397
H@1 0.6560 0.2640 0.7080 0.6640 0.6480 0.6520 0.6600 0.6640
P@5 0.6144 0.2456 0.6168 0.6176 0.6184 0.5968 0.6160 0.6240
Amazon Books Reviews N@10 0.9085 0.3134 0.9831 0.9860 0.9842 0.9695 0.9868 0.9883
H@1 0.8305 0.2800 0.9840 0.9880 0.9760 0.9400 0.9840 0.9960
P@5 0.7720 0.2632 0.9320 0.9352 0.9336 0.9192 0.9416 0.9400
MIND N@10 0.8414 0.6435 0.8372 0.8351 0.8426 0.8374 0.8430 0.8368
H@1 0.6920 0.4800 0.6480 0.6480 0.6840 0.6720 0.6880 0.6640
P@5 0.6096 0.4412 0.6256 0.6208 0.6256 0.6208 0.6208 0.6152
Anime Dataset 2023 N@10 0.9733 0.2918 0.9710 0.9655 0.9693 0.9695 0.9731 0.9717
H@1 0.9480 0.2960 0.9480 0.9200 0.9440 0.9400 0.9560 0.9440
P@5 0.9304 0.2528 0.9208 0.9152 0.9144 0.9192 0.9264 0.9200
Cross-Domain Recommendation
Amazon Reviews 2023 N@10 0.9018 0.8859 0.8015 0.9037 0.8984 0.9036 0.8928 0.9026
H@1 0.8144 0.8084 0.7305 0.8563 0.8024 0.8323 0.7725 0.8323
P@5 0.7281 0.6946 0.6347 0.7353 0.7341 0.7485 0.7257 0.7413
Next Item Recommendation
Amazon Reviews 2023 H@1 0.7965 0.8268 0.8182 0.5671 0.7879 0.8095 0.8268 0.7922
MicroLens H@1 0.9560 0.9440 0.9400 0.4360 0.9400 0.9560 0.9320 0.9040
MovieLens 32M H@1 0.6280 0.6040 0.6000 0.4480 0.6360 0.5840 0.6040 0.6040
PixelRec H@1 0.8520 0.8880 0.8640 0.2360 0.8760 0.8480 0.8480 0.8800
  • *“RP” stands for “Rating Prediction”, “TKR” means “Top-K Recommendation”, “CDR” means “Cross-Domain Recommendation”, “NIR” refers to “Next Item Recommendation”, “UAP” indicates “User Attribute Prediction”, “IR” indicates “Interest Recognition”, “TUI” denotes “Target User Identification”, and ”TD” represents ”Task Description”.
Table 13. Results for user-item understanding tasks on GLM-4-9B with ITDR removing different subtasks.
Task & Dataset Metric w/o RP w/o TKR w/o CDR w/o NIR w/o UAP w/o IR w/o TUI w/o TD
User Attribute Prediction
Last.FM 1K ACC(country) 0.5405 0.6486 0.5676 0.6216 0.2432 0.5676 0.4865 0.5946
ACC(gender) 0.4595 0.6216 0.6216 0.4865 0.3243 0.5946 0.5135 0.5135
RMSE(age) 6.4512 6.5533 6.8569 6.1045 8.2658 6.0589 6.1298 6.8821
Last.FM 360K ACC(country) 0.4600 0.4720 0.4480 0.4320 0.2240 0.4360 0.4680 0.4800
ACC(gender) 0.7520 0.7240 0.7520 0.7400 0.6600 0.7400 0.7320 0.7160
RMSE(age) 8.9140 9.0841 8.6457 8.9904 10.1927 8.4669 8.8367 8.7964
Anime Dataset 2023 ACC(location) 0.0320 0.0480 0.0600 0.0360 0.0080 0.0440 0.0400 0.0400
ACC(gender) 0.8120 0.8080 0.8240 0.7960 0.6560 0.8160 0.8440 0.8320
RMSE(Birth year) 4.6152 4.6673 4.3474 4.5760 6.8148 4.6411 4.4231 4.3433
BookCrossing ACC(location) 0.0560 0.0720 0.0600 0.0400 0.0000 0.0520 0.0680 0.0680
RMSE(age) 16.5817 15.6690 16.4448 15.5870 18.2569 15.6729 15.5358 15.1733
MovieLens 1M ACC(age group) 0.3750 0.3958 0.3583 0.3875 0.3667 0.3750 0.4000 0.3708
ACC(gender) 0.7542 0.7792 0.7417 0.7750 0.5375 0.7250 0.7458 0.7375
ACC(occupation) 0.1292 0.1458 0.1500 0.1292 0.0000 0.1542 0.1168 0.0958
Interest Recognition
Amazon Books Reviews BLEU-1 0.3571 0.3502 0.3446 0.3515 0.3481 0.1782 0.3549 0.3604
BLEU-2 0.2235 0.2174 0.2156 0.2174 0.2164 0.0865 0.2177 0.2255
ROUGE-1 0.4185 0.4138 0.4119 0.4156 0.4135 0.2782 0.4135 0.4210
ROUGE-2 0.1811 0.1754 0.1753 0.1753 0.1754 0.0758 0.1731 0.1799
ROUGE-L 0.3941 0.3890 0.3865 0.3899 0.3890 0.2598 0.3897 0.3944
Anime Dataset 2023 BLEU-1 0.4424 0.4449 0.4455 0.4486 0.4386 0.2371 0.4394 0.4379
BLEU-2 0.3337 0.3362 0.3371 0.3336 0.3290 0.1230 0.3316 0.3308
ROUGE-1 0.5435 0.5440 0.5400 0.5455 0.5350 0.3120 0.5388 0.5409
ROUGE-2 0.3040 0.3056 0.3053 0.2982 0.2965 0.0857 0.3003 0.3025
ROUGE-L 0.5084 0.5080 0.5081 0.5101 0.4996 0.2816 0.5046 0.5057
Steam BLEU-1 0.4676 0.4626 0.4626 0.4670 0.4730 0.2126 0.4582 0.4631
BLEU-2 0.3166 0.3131 0.3128 0.3145 0.3221 0.1208 0.3091 0.3123
ROUGE-1 0.5000 0.4945 0.4900 0.4955 0.4993 0.3482 0.4928 0.4964
ROUGE-2 0.2395 0.2365 0.2345 0.2347 0.2422 0.1117 0.2332 0.2359
ROUGE-L 0.4669 0.4618 0.4586 0.4614 0.4666 0.3183 0.4590 0.4623
  • *“RP” stands for “Rating Prediction”, “TKR” means “Top-K Recommendation”, “CDR” means “Cross-Domain Recommendation”, “NIR” refers to “Next Item Recommendation”, “UAP” indicates “User Attribute Prediction”, “IR” indicates “Interest Recognition”, “TUI” denotes “Target User Identification”, and ”TD” represents ”Task Description”.
Table 14. Continued from Table 13.
Task & Dataset Metric w/o RP w/o TKR w/o CDR w/o NIR w/o UAP w/o IR w/o TUI w/o TD
Target User Identification
Amazon Reviews 2023 BLEU-1 0.4892 0.4858 0.4920 0.4856 0.4909 0.4827 0.0363 0.4866
BLEU-2 0.3381 0.3328 0.3387 0.3329 0.3382 0.3320 0.0184 0.3353
ROUGE-1 0.5203 0.5138 0.5182 0.5116 0.5165 0.5137 0.2261 0.5172
ROUGE-2 0.2603 0.2536 0.2588 0.2530 0.2575 0.2552 0.0547 0.2580
ROUGE-L 0.4899 0.4856 0.4887 0.4834 0.4875 0.4859 0.1995 0.4870
BookCrossing BLEU-1 0.5088 0.5125 0.5052 0.5090 0.5132 0.5096 0.0073 0.5072
BLEU-2 0.3605 0.3601 0.3552 0.3576 0.3639 0.3580 0.0039 0.3582
ROUGE-1 0.5377 0.5429 0.5363 0.5386 0.5454 0.5412 0.1663 0.5384
ROUGE-2 0.2783 0.2763 0.2734 0.2744 0.2830 0.2747 0.0419 0.2771
ROUGE-L 0.5063 0.5100 0.5053 0.5062 0.5143 0.5091 0.1467 0.5070
Anime Dataset 2023 BLEU-1 0.5463 0.5469 0.5497 0.5461 0.5520 0.5452 0.0064 0.5492
BLEU-2 0.4074 0.4067 0.4065 0.4085 0.4110 0.4040 0.0037 0.4098
ROUGE-1 0.5809 0.5831 0.5845 0.5841 0.5885 0.5825 0.1515 0.5859
ROUGE-2 0.3275 0.3270 0.3251 0.3324 0.3306 0.3244 0.0344 0.3309
ROUGE-L 0.5502 0.5497 0.5516 0.5534 0.5572 0.5508 0.1315 0.5544
Steam BLEU-1 0.5104 0.5144 0.5090 0.5144 0.5158 0.5125 0.0942 0.5121
BLEU-2 0.3686 0.3711 0.3641 0.3699 0.3706 0.3703 0.0509 0.3697
ROUGE-1 0.5389 0.5401 0.5355 0.5409 0.5367 0.5389 0.2977 0.5396
ROUGE-2 0.2866 0.2868 0.2787 0.2852 0.2833 0.2866 0.0857 0.2865
ROUGE-L 0.5087 0.5101 0.5044 0.5086 0.5067 0.5095 0.2597 0.5097
Yelp BLEU-1 0.4824 0.4819 0.4872 0.4826 0.4842 0.4848 0.0786 0.4849
BLEU-2 0.3292 0.3312 0.3325 0.3293 0.3327 0.3310 0.0404 0.3318
ROUGE-1 0.5149 0.5185 0.5182 0.5188 0.5185 0.5184 0.2552 0.5192
ROUGE-2 0.2523 0.2575 0.2534 0.2541 0.2572 0.2540 0.0676 0.2563
ROUGE-L 0.4870 0.4913 0.4899 0.4896 0.4903 0.4913 0.2299 0.4901
MicroLens BLEU-1 0.4996 0.5023 0.4976 0.5047 0.5011 0.5001 0.0025 0.5049
BLEU-2 0.3581 0.3617 0.3560 0.3640 0.3614 0.3600 0.0010 0.3628
ROUGE-1 0.5331 0.5360 0.5317 0.5399 0.5396 0.5369 0.1295 0.5393
ROUGE-2 0.2809 0.2850 0.2797 0.2871 0.2872 0.2843 0.0272 0.2848
ROUGE-L 0.5036 0.5082 0.5026 0.5103 0.5109 0.5094 0.1147 0.5086
PixelRec BLEU-1 0.4843 0.4871 0.4869 0.4874 0.4866 0.4849 0.0012 0.4896
BLEU-2 0.3436 0.3445 0.3452 0.3453 0.3437 0.3438 0.0006 0.3486
ROUGE-1 0.5203 0.5237 0.5221 0.5242 0.5226 0.5222 0.1243 0.5271
ROUGE-2 0.2656 0.2663 0.2676 0.2676 0.2649 0.2665 0.0281 0.2707
ROUGE-L 0.4885 0.4926 0.4920 0.4921 0.4904 0.4912 0.1093 0.4955
MovieLens 32M BLEU-1 0.5311 0.5232 0.5268 0.5210 0.5227 0.5196 0.0021 0.5247
BLEU-2 0.3876 0.3806 0.3830 0.3758 0.3796 0.3763 0.0013 0.3806
ROUGE-1 0.5529 0.5482 0.5504 0.5456 0.5493 0.5454 0.1375 0.5502
ROUGE-2 0.2997 0.2950 0.2956 0.2890 0.2945 0.2915 0.0331 0.2940
ROUGE-L 0.5154 0.5102 0.5116 0.5054 0.5109 0.5081 0.1186 0.5111
MIND BLEU-1 0.5578 0.5571 0.5574 0.5587 0.5520 0.5534 0.0010 0.5627
BLEU-2 0.4034 0.4029 0.3993 0.4043 0.3972 0.4004 0.0006 0.4072
ROUGE-1 0.6104 0.6098 0.6097 0.6096 0.6084 0.6108 0.1339 0.6153
ROUGE-2 0.3194 0.3188 0.3131 0.3198 0.3151 0.3199 0.0351 0.3233
ROUGE-L 0.5777 0.5773 0.5752 0.5780 0.5754 0.5789 0.1215 0.5824
Table 15. Experimental results of closed source LLMs on user-item interaction tasks.
Task & Dataset Metric GPT-3.5-Turbo Gemini 2.5 Flash
Rating Prediction
MovieLens 32M RMSE 0.9775 0.9644
BookCrossing RMSE 1.6583 1.8947
Anime Dataset 2023 RMSE 1.4000 1.4156
Yelp RMSE 1.5602 2.1422
Top-K Recommendation
Amazon Reviews 2023 N@10 0.8630 0.8682
H@1 0.7797 0.8093
P@5 0.6576 0.6941
MovieLens 32M N@10 0.9696 0.9685
H@1 0.9800 0.9440
P@5 0.8920 0.8856
MicroLens N@10 0.7854 0.8009
H@1 0.6760 0.7200
P@5 0.5528 0.6000
PixelRec N@10 0.7241 0.7687
H@1 0.5840 0.7000
P@5 0.4752 0.5456
Amazon Books Reviews N@10 0.9661 0.9239
H@1 0.9600 0.9320
P@5 0.8080 0.8376
MIND N@10 0.7598 0.7523
H@1 0.5960 0.6440
P@5 0.5237 0.5304
Anime Dataset 2023 N@10 0.9790 0.9527
H@1 0.9920 0.9200
P@5 0.8808 0.8616
Cross-Domain Recommendation
Amazon Reviews 2023 N@10 0.8111 0.8335
H@1 0.6048 0.7605
P@5 0.5928 0.6419
Next Item Recommendation
Amazon Reviews 2023 H@1 0.3723 0.5801
MicroLens H@1 0.3440 0.5280
MovieLens 32M H@1 0.3960 0.4360
PixelRec H@1 0.2880 0.2400
Table 16. Experimental results of closed source LLMs on user-item understanding tasks.
Task & Dataset Metric GPT-3.5-Turbo Gemini 2.5 Flash
User Attribute Prediction
Last.FM 1K ACC(country) 0.1622 0.3514
ACC(gender) 0.5405 0.6757
RMSE(age) 5.0929 7.7127
Last.FM 360K ACC(country) 0.2440 0.4080
ACC(gender) 0.5360 0.5280
RMSE(age) 9.1007 11.2361
Anime Dataset 2023 ACC(location) 0.0240 0.0000
ACC(gender) 0.5560 0.6840
RMSE(Birth year) 5.4794 5.5656
BookCrossing ACC(location) 0.0800 0.0800
RMSE(age) 14.5382 15.6048
MovieLens 1M ACC(age group) 0.3167 0.0667
ACC(gender) 0.7208 0.7250
ACC(occupation) 0.0083 0.0167
Interest Recognition
Amazon Books Reviews BLEU-1 0.1551 0.1030
BLEU-2 0.0675 0.0386
ROUGE-1 0.2098 0.1706
ROUGE-2 0.0420 0.0245
ROUGE-L 0.1368 0.0975
Anime Dataset 2023 BLEU-1 0.1452 0.1412
BLEU-2 0.0830 0.0559
ROUGE-1 0.2560 0.2032
ROUGE-2 0.0771 0.0332
ROUGE-L 0.1662 0.1497
Steam BLEU-1 0.1753 0.1014
BLEU-2 0.1046 0.0469
ROUGE-1 0.3075 0.2297
ROUGE-2 0.1177 0.0523
ROUGE-L 0.2032 0.1335
Table 17. Continued from Table 16.
Task & Dataset Metric GPT-3.5-Turbo Gemini 2.5 Flash
Target User Identification
Amazon Reviews 2023 BLEU-1 0.3182 0.2862
BLEU-2 0.1811 0.1474
ROUGE-1 0.3644 0.3277
ROUGE-2 0.1275 0.0946
ROUGE-L 0.2229 0.1854
BookCrossing BLEU-1 0.2481 0.2751
BLEU-2 0.1512 0.1455
ROUGE-1 0.3786 0.3436
ROUGE-2 0.1433 0.1013
ROUGE-L 0.2323 0.2008
Anime Dataset 2023 BLEU-1 0.2983 0.2499
BLEU-2 0.1774 0.1303
ROUGE-1 0.3799 0.3267
ROUGE-2 0.1382 0.0922
ROUGE-L 0.2311 0.1886
Steam BLEU-1 0.3050 0.2267
BLEU-2 0.1783 0.1130
ROUGE-1 0.3703 0.3027
ROUGE-2 0.1322 0.0779
ROUGE-L 0.2241 0.1713
Yelp BLEU-1 0.3122 0.2694
BLEU-2 0.1784 0.1419
ROUGE-1 0.3622 0.3278
ROUGE-2 0.1238 0.0960
ROUGE-L 0.2113 0.1914
MicroLens BLEU-1 0.1700 0.3419
BLEU-2 0.1185 0.1996
ROUGE-1 0.3758 0.3722
ROUGE-2 0.1806 0.1336
ROUGE-L 0.2569 0.2299
PixelRec BLEU-1 0.1893 0.3338
BLEU-2 0.1159 0.1938
ROUGE-1 0.3416 0.3675
ROUGE-2 0.1303 0.1304
ROUGE-L 0.2093 0.2248
MovieLens 32M BLEU-1 0.2281 0.2626
BLEU-2 0.1358 0.1456
ROUGE-1 0.3478 0.3382
ROUGE-2 0.1288 0.1098
ROUGE-L 0.2079 0.2096
MIND BLEU-1 0.2839 0.3327
BLEU-2 0.1705 0.1875
ROUGE-1 0.4134 0.4093
ROUGE-2 0.1455 0.1297
ROUGE-L 0.2378 0.2436
Table 18. A data example of the rating prediction task. Data source: Anime Dataset 2023.
Rating Prediction
Task Description:
The objective of the rating prediction task is to thoroughly explore users’ preferences toward items (e.g., movies, products, music, books, etc.) based on their historical rating records. By extracting insights from the rating data to identify the types of items users prefer or dislike, the task aims to accurately predict users’ ratings for new items.
Input:
A user has watched and rated the following anime with ratings: Bleach (7), Naruto: Shippuuden (6), Bishoujo Senshi Sailor Moon (7), Bishoujo Senshi Sailor Moon S: Kaguya-hime no Koibito (4), Bishoujo Senshi Sailor Moon SuperS (4), Bishoujo Senshi Sailor Moon SuperS Specials (4), Bleach Movie 1: Memories of Nobody (7), Bleach Movie 3: Fade to Black - Kimi no Na wo Yobu (6), Bleach: Memories in the Rain (6), Bleach: The Sealed Sword Frenzy (6). Predict his rating for Chrno Crusade.
Output:
9
Table 19. A data example of the top k recommendation task. Data source: PixelRec.
Top-K Recommendation
Task Description:
The primary goal of Top-K recommendation is to select the K items most likely to attract the user from a candidate set based on their historical behavior or preferences, and present them in a ranked list.
Input:
Given a user’s short video watching records in chronological order as follows:
1. How to get the most out of your favorite vegetables in the world? Why are all American vegetables the same?
2. Qingming Festival in Xi’an storm storm eating, get up first drink eight bowls of wrestling bowl wine!
3. The first time I made bean flowers in the countryside, Wang Gang asked me to challenge the spiciest seven-star pepper!
4. Do American college students look for their families for living expenses and tuition? Street interviews with Americans!
5. [European and American group portrait] ”You are the hopeless lover I hide under the great friendship.”
Based on the watching history, please rank the short video in the candidate list by the likelihood of the user watching them next from highest to lowest.
Candidate list:
1. Game of Thrones Male Character Homo Popularity Rankings
2. The fat woman at home, forcing her pregnant daughter to divorce to take care of herself, stealing food to lose weight and scolding the doctor!
3. Free to receive the original price of Mars Survival Deluxe Edition, come to white whoring hi plus one!
4. Chemistry ”Stalactites”
5. [Emma Slippery Won] Facing the straight ball sweet girl, he is really panicked!
6. All up, you must know this niche and good treasure app!
7. Zeta Ultraman meets all generations of Ultraman
8. Game of Thrones: How many times did Haima Raijin laugh at the most hepatic collection on the B station?
9. I, Killsumaru, have no one to protect.
10. President Seahorse
Output:
1. Game of Thrones: How many times did Haima Raijin laugh at the most hepatic collection on the B station?
2. Game of Thrones Male Character Homo Popularity Rankings
3. President Seahorse
4. I, Killsumaru, have no one to protect.
5. Zeta Ultraman meets all generations of Ultraman
6. The fat woman at home, forcing her pregnant daughter to divorce to take care of herself, stealing food to lose weight and scolding the doctor!
7. [Emma Slippery Won] Facing the straight ball sweet girl, he is really panicked!
8. Free to receive the original price of Mars Survival Deluxe Edition, come to white whoring hi plus one!
9. All up, you must know this niche and good treasure app!
10. Chemistry ”Stalactites”
Table 20. A data example of the cross-domain recommendation task. Data source: Amazon Review 2023.
Cross-Domain Recommendation
Task Description:
The goal of cross-domain recommendation is to predict a user’s behavior in the target domain based on their historical behavior in the source domain, thereby achieving effective utilization of cross-domain information and improving recommendation performance.
Input:
A user’s purchase history in the beauty domain is as follows:
1. FIXBODY Boar Bristle Hair Brush (Wooden)
2. Hair Scrunchies for Women & Girls , 25PCS Hair Scrunchies Accessories Included Velvet, Chiffon, Flower Printed, Silk Hair Ties Band & Hair Bun Maker, Elastic Hair Scrunchies Perfect for Hair Bun, Ponytail Holder
3. Liquid Foundation Cream, Concealer Cover Cream, BB Cream, Flawless Colour Changing Foundation Makeup Base Nude Face Liquid Cover Concealer, Universal for ALL Skin Types, SPF 15 Shell, 1.41 Fl Oz
4. Liquid Foundation,Foundation Cream,Flawless Colour Changing Foundation,Hides Wrinkles & Lines,BB Cream,Covering Imperfections Liquid Complete Foundation Cover,Universal Shade for ALL Skin Makeup
5. Morovan 8 Colors Gel Nail Polish Kit with 48W UV LED Light Nail Lamp Pink White Blue Purple Colors For Women With Base and Top Coat Full DIY Professional Gel Manicure Nail Tools for Home and Salon Use,
and in the fashion domain is as follows: 1. MITALOO Push Up Strapless Sticky Adhesive Invisible Backless Bras Plunge Reusable Magic Bra for Women Beige
2. ENIDMIL Womens Soft Cotton Pajama Set Short Sleeve Top and Long Pants Bottom Sleepwear Loungewear PJ Set Pajamas for Women (Viola Purple, Large)
3. Fettero Women Simple 14K Gold Plated Chain Delicate Single Freshwater Pearls Double Layer Choker Necklace
4. Rave Hood 3D Bandana Gaiter Headwear Headband Magic Scarf Seamless 2019 Improved for Fishing, Man Women Running, Hiking, Motorcycle
5. Three Dots Pendants Necklaces,14k Gold Plated 3 Dot Circle Disks Chain Tiny Triple Round Flat Bead Dainty Handmade Pendants Necklaces Jewelry Gifts for Women Girls.
Based on this information, please rank the following list of products in the fashion domain in order of the likelihood that the user will purchase them next, from highest to lowest. Product list:
1. BSJELL Straw Earrings Woven Handmade Rattan Geometric Drop Earrings Hammered Disc Stud Wicker Bohemian Lightweight Earrings for Women (Teardop)
2. BIADANI Women Versatile Long Sleeve Open Front Drape Front Pockets Cardigan Hacci Pink XX-Large
3. Stainless Steel Men Cuff Link Cuff Links Shirt Set Tile
4. Tassel Fringe Earrings Drop Dangle Earring for Women Girls Black by THSjewelry
5. TOBWIZU Sports Bras for Women,Seamless Sleep Bra with Removable Pads Yoga Gym Activity Everyday Wear
6. DEDU MusBema Kids Casual Lightweight Breathable Sneakers Easy Walk Sport Shoes for Boys Girls White
7. Wannpoty Women’s Summer Dress Sleeveless Casual Loose Button T-Shirt with Double Pockets (Large, Dark Green)
8. Oriental Pearl Mama Shirts Women Girl Mom Letter Printed T-Shirt Mama Graphic Tee Casual Short Sleeve Tops Tee (Pink2, XXL)
9. GemsChest 925 Sterling Silver Ring for Women Round Cubic Zirconia 3-Stone Engagement Wedding Jewelry Hola CZ Bridal Ring
10. black Crossbody bags for women Lightweight Medium Classic Bag with tassel and Zipper Pocket Adjustable Strap PU leather.
Output:
1. BSJELL Straw Earrings Woven Handmade Rattan Geometric Drop Earrings Hammered Disc Stud Wicker Bohemian Lightweight Earrings for Women (Teardop)
2. Tassel Fringe Earrings Drop Dangle Earring for Women Girls Black by THSjewelry
3. black Crossbody bags for women Lightweight Medium Classic Bag with tassel and Zipper Pocket Adjustable Strap PU leather
4. Wannpoty Women’s Summer Dress Sleeveless Casual Loose Button T-Shirt with Double Pockets (Large, Dark Green)
5. GemsChest 925 Sterling Silver Ring for Women Round Cubic Zirconia 3-Stone Engagement Wedding Jewelry Hola CZ Bridal Ring
6. TOBWIZU Sports Bras for Women,Seamless Sleep Bra with Removable Pads Yoga Gym Activity Everyday Wear
7. DEDU MusBema Kids Casual Lightweight Breathable Sneakers Easy Walk Sport Shoes for Boys Girls White
8. BIADANI Women Versatile Long Sleeve Open Front Drape Front Pockets Cardigan Hacci Pink XX-Large
9. Oriental Pearl Mama Shirts Women Girl Mom Letter Printed T-Shirt Mama Graphic Tee Casual Short Sleeve Tops Tee (Pink2, XXL)
10. Stainless Steel Men Cuff Link Cuff Links Shirt Set Tile
Table 21. A data example of the next item recommendation task. Data source: MicroLens.
Next Item Recommendation
Task Description:
The primary goal of next-item recommendation is to predict the item that a user is most likely to interact with or be interested in next, based on their historical behavior sequence (such as clicks, purchases, or views).
Input:
Given that a short video user has watched the following videos in chronological order: [’1.planted sweet potato seedlings after the rain:Delicacy’, ’2.the field of cowpeas fast old:Delicacy’, ’3.under the blue sky and white clouds small Viagra with Xiang sister to touch the crunch:Delicacy’, ’4.far married daughter-in-law pregnant want to eat chicken:Daily Sharing’, ’5.3 years old cute baby want to super restaurant:Daily Sharing’], please select the next video the user is most likely to watch from the candidate list: [’(follow-up) Yesterday:Daily Sharing’, ’Let’s go! # Bear-infested # quadratic nova:Anime’, ’the sixth set of Diorman princess:Film and Television Works’, ’in order to make snacks for her son-in-law:Daily Sharing’, ’Hongli for the rescue of the green Ying does not put the empress dowager access to Cining Palace:Film and Television Works’]
Output:
in order to make snacks for her son-in-law
Table 22. A data example of the user attribute prediction task. Data source: MovieLens 1M.
User Attribute Prediction
Task Description:
The objective of attribute prediction is to infer implicitly expressed static attributes (e.g., age, gender, income, education level) from user behavioral data (e.g., browsing, purchases, social interactions), thereby completing fundamental dimensions of user profiles to support differentiated recommendation strategies.
Input:
The user’s movie preferences are as follows:
He/She likes watching Dead Man Walking (1995): Drama, Dead Poets Society (1989): Drama, Drugstore Cowboy (1989): Crime|Drama, The Graduate (1967): Drama|Romance, North by Northwest (1959): Drama|Thriller, The Insider (1999): Drama, 12 Angry Men (1957): Drama, A Christmas Story (1983): Comedy|Drama, On the Waterfront (1954): Crime|Drama, The Shawshank Redemption (1994): Drama, The Silence of the Lambs (1991): Drama|Thriller, To Kill a Mockingbird (1962): Drama, The Boat(Das Boot) (1981): Action|Drama|War, 2001: A Space Odyssey (1968): Drama|Mystery|Sci-Fi|Thriller, Midnight Cowboy (1969): Drama, Serpico (1973): Crime|Drama, The Wizard of Oz (1939): Adventure|Children’s|Drama|Musical, American Beauty (1999): Comedy|Drama, Citizen Kane (1941): Drama, Casablanca (1942): Drama|Romance|War.
Based on the user’s movie preferences, predict his/her age group, gender, and occupation.
Output:
age group:25-34,gender:M,occupation:doctor/health care.
Table 23. A data example of the interest recognition task. Data source: Amazon Books Reviews.
Interest Recognition
Task Description:
The objective of interest recognition is to infer and categorize user interests based on behavioral data, thereby providing interpretable and actionable interest dimensions for personalized recommendation systems.
Input:
A user has given the following book reviews:
Liked books:
Little House on the Prairie: I so love these books…I am in my 30s and and just now reading them and they are such a delight…as a child i loved the tv series and the books I must say are even better…
Little Lord Fauntleroy: my goodness this is one of the best books ive read….which i am a BIG fan of fhb and have liked everything i have read written by her…this has to be one of my favorites….great story great meaning to the story!!! please read!!!
My Sweet Audrina: This was a pretty good read thought I figured out the s̈ecretm̈idway through the book. It’s really kind of obvious if you think about it!!! overall thought this was a good book.
Disliked books:
Silas Marner: The weaver of Raveloe, (The Academy series of English classics): This has to be the most boring book i have ever read! It seems like it takes like 10 pages just to say what could be said in one paragraph. oh my goodness completely boring please dont put your self through this.
Can’t Wait to Get to Heaven: A Novel: This book was SO disappointing to me. I am a Christin woman and don’t like to read accounts of Heaven not taken from the Bible. I first started this book and thought it was going to be great. I stopped on page 134 couldn’t go any further. The story was dragged out and overall boring. Waste of time!
Silas Marner (New Windmill): This has to be the most boring book i have ever read! It seems like it takes like 10 pages just to say what could be said in one paragraph. oh my goodness completely boring please dont put your self through this.
Based on these reviews, briefly summarize the user’s reading interests, focusing on the types of books, styles, and authors the user may prefer.
Output:
The user enjoys classic, wholesome, and sentimental stories (e.g., Little House on the Prairie, Little Lord Fauntleroy) with strong moral themes and emotional depth, possibly favoring older or traditional authors (Frances Hodgson Burnett). They dislike slow-paced, overly verbose classics (Silas Marner) and non-biblical religious fiction (Can’t Wait to Get to Heaven). They also appreciate mystery/drama with twists (*My Sweet Audrina) but prefer clear, engaging storytelling without excessive filler.In short: Classic, heartfelt, morally uplifting books with brisk pacing; dislikes tedious or theologically speculative works.
Table 24. A data example of the target user identification task. Data source: Steam.
Target User Identification
Task Description:
The objective of target user identification is to accurately pinpoint the user groups most likely to exhibit positive feedback (e.g., clicks, purchases, high ratings) by analyzing the alignment between item characteristics and user profiles, thereby providing targeted distribution guidance for recommendation systems.
Input:
The attributes of the game Cooking Academy Fire and Knives are as follows:
genres: [’Action’, ’Casual’],
tags: [’Casual’, ’Action’, ’Simulation’],
specs: [’Single-player’], publisher: Viva Media,
Customers’ comments on the game are as follows:
1. Just theraputic & fun to play if you enjoy cooking.
2. 10/10 I am the next Rachel Ray now because this game taught me how to crack an egg properly.
3. Fun and very nice to play. Can be a bit stressful at times though.
Based on the attributes and the reviews, briefly describe which user groups this game is suitable for recommending.
Output:
Based on the game’s attributes and customer reviews, Cooking Academy Fire and Knives is best suited for the following user groups:
1. Casual Gamers – The game’s genres (Action, Casual) and tags (Casual, Simulation) suggest it’s easy to pick up and play without requiring deep commitment.
2. Cooking Enthusiasts – Reviews highlight its therapeutic and educational aspects (e.g., learning cooking techniques like cracking eggs), making it appealing to those who enjoy cooking-themed games.
3. Players Seeking Relaxing Yet Engaging Experiences – While mostly fun and therapeutic, the slight stress mentioned in reviews suggests it offers light challenges, fitting players who enjoy mild intensity.
4. Single-Player Gamers – Since it’s single-player only, it suits those who prefer solo gameplay.
Avoid recommending to competitive multiplayer fans or those seeking hardcore action/simulation depth.